### v2

# TSD-SR: One-Step Diffusion with Target Score Distillation for Real-World Image Super-Resolution

Abstract
Pre-trained text-to-image diffusion models are increasingly applied to real-world image super-resolution (Real-ISR) task. Given the iterative refinement nature of diffusion models, most existing approaches are computationally expensive. While methods such as SinSR and OSEDiff have emerged to condense inference steps via distillation, their performance in image restoration or details recovery is not satisfied. To address this, we propose TSD-SR, a novel distillation framework specifically designed for real-world image super-resolution, aiming to construct an efficient and effective one-step model. We first introduce the Target Score Distillation, which leverages the priors of diffusion models and real image references to achieve more realistic image restoration. Secondly, we propose a Distribution-Aware Sampling Module to make detail-oriented gradients more readily accessible, addressing the challenge of recovering fine details. Extensive experiments demonstrate that our TSD-SR has superior restoration results (most of the metrics perform the best) and the fastest inference speed (e.g. 40 times faster than SeeSR) compared to the past Real-ISR approaches based on pre-trained diffusion priors. Our code will be released at https://github.com/Microtreei/TSD-SR.
1 Introduction
Image super-resolution (ISR) [9, 8, 23, 26] aims to transform low-quality (LQ) images, which have experienced noise or blur, into clear high-quality (HQ) images. Differing from traditional ISR [6, 66], which assumes a known degradation process, real-world image super-resolution (Real-ISR) [46, 62] is designed to enhance real-world images that have suffered from complex and unknown degradations, thereby offering greater practical utility.
Generative models, particularly Generative Adversarial Networks (GANs) [11, 33, 37] and Diffusion Models (DMs) [41, 17, 39], have shown remarkable power in handling Real-ISR task. GAN-based methods leverage adversarial training, toggling between the generator and discriminator to produce realistic images. While GANs are capable of one-step inference, they are commonly hampered by issues like mode collapse and training instability [2]. Recently, Diffusion Models (DMs) have shown impressive performance in the realm of image generation [47, 21]. Their robust priors empower them to produce more realistic images with richer details than GAN-based methods [39, 41]. Some researchers [29, 60, 53, 56] have successfully leveraged pre-trained DMs for Real-ISR task. However, due to the iterative denoising nature of diffusion models [17], the Real-ISR process is computationally expensive.
Figure 1: Performance and efficiency comparison among Real-ISR methods. TSD-SR stands out for achieving high-quality restoration with the fastest speed among diffusion-based models. In contrast, existing models prioritize either speed or restoration performance. The performance of each method is benchmarked on an A100 GPU with the DRealSR dataset.
To achieve an efficient and one-step network akin to GANs, several pioneering methods that condense the iterations of diffusion models through distillation [15, 57, 12, 18] have been proposed [52, 54, 48]. Among these works, OSEDiff [52] introduced the Variational Score Distillation (VSD) loss [50] to Real-ISR task, achieving state-of-the-art (SOTA) one-step performance by leveraging prior knowledge from pre-trained models. Despite these advancements, our investigation has revealed two critical limitations associated with VSD in Real-ISR applications. (1) Unreliable gradient direction. VSD relies on a Teacher Model to provide a “true gradient direction.” However, this guidance proves unreliable in scenarios where initial ISR outputs are suboptimal. (2) Insufficient detail recovery. The VSD loss exhibits notable variation across different timesteps, and the uniform sampling strategy for $t$ poses challenges in aligning the score function with detailed texture recovery requirements. These findings highlight the need for enhanced approaches to address these issues effectively.
In this paper, we propose a novel method called TSD-SR to distill a multi-step Text-to-Image (T2I) DMs [38, 39, 10] into an effective one-step diffusion model tailored for the Real-ISR task. Specifically, TSD-SR consists of two components: Target Score Distillation (TSD) and Distribution-Aware Sampling Module (DASM). TSD incorporates our new proposed Target Score Matching (TSM) loss to compensate for the limitation of VSD loss. This significant score loss leverages HQ data to provide a reliable optimization trajectory for the distillation process. It effectively reduces the visual artifacts caused by deviant predictions from the Teacher Model. DASM is crafted to bolster detail recovery by strategically sampling low-noise samples that are distribution-based during training. This approach effectively allocates more optimization to early timesteps in a single iteration, enhancing the recovery of details.
Experiments on popular benchmarks demonstrate that TSD-SR achieves superior restoration performance (most of the metrics perform the best) and high efficiency (the fastest inference speed, 40 times faster than SeeSR) compared to the state-of-the-art Real-ISR methods based on pre-trained DMs, while requiring only a single inference step.
Our main contribution can be summarized as threefold:
  • We propose a novel method called TSD-SR to achieve one-step DMs distillation for the Real-ISR task.
  • We introduce Target Score Distillation (TSD) to provide a reliable gradient to enhance the realism of the Real-ISR model’s outputs.
  • We design a Distribution-Aware Sampling Module (DASM) specifically tailored to enhance the capability of detail restoration.
2 Related Work
GAN-based Real-ISR. Since SRGAN [26] first applied GAN to ISR, it has effectively enhanced visual quality by combining adversarial loss with perceptual loss [65, 7]. Subsequently, ESRGAN [45] introduced Residual-in-Residual Dense Block and a relativistic average discriminator, further improving detail restoration. Methods like BSRGAN [62] and Real-ESRGAN [46] simulate complex real-world degradation processes, achieving ISR under unknown degradation conditions, which enhances the model’s generalization ability. Although GAN-based methods are capable of adding more realistic details to images, they suffer from training instability and mode collapse [2].
Multi-step Diffusion-based Real-ISR. Some researches [44, 29, 53, 60, 56] in recent years have utilized the powerful image priors in pre-trained T2I diffusion models [64, 35, 39] for Real-SR tasks and achieved promising results. For example, StableSR [44] balances fidelity and perceptual quality by fine-tuning the time-aware encoder and employing controllable feature wrapping. DiffBiR [29] first processes the LR image through a reconstruction network and then uses the Stable Diffusion (SD) model [39] to supplement the details. SeeSR [53] attempts to better stimulate the generative power of the SD model by extracting the semantic information in the image as a conditional guide. PASD [56] introduces a pixel-aware cross attention module to enable the diffusion model to perceive the local structure of the image at the pixel level, while using a degradation removal module to extract degradation insensitive features to guide the diffusion process along with high-level information from the image. SUPIR [60] achieves a generative and fidelity capability using negative cues [16] as well as restoration-guided sampling, while using a larger pre-training model with a larger dataset to enhance the model capability. However, all of these methods are limited by the multi-step denoising of the diffusion model, which requires 20-50 iterations in inference, resulting in an inference time that lags far behind that of GAN-based methods.
One-step Diffusion-based Real-ISR. Recently, there has been a surge of interest within the academic community in one-step distillation techniques [34, 59, 40, 58, 31] for diffusion-based Real-ISR task. SinSR [48] leverages consistency preserving distillation to condense the inference steps of ResShift [61] into a single step, yet the generalization of ResShift and SinSR is constrained due to the absence of large-scale data training. AddSR [54] introduces the adversarial diffusion distillation (ADD) [40] to Real-ISR tasks, resulting in a comparatively effective four-step model. However, this method has a propensity to produce excessive and unnatural image details. OSEDiff [52] directly uses LQ images as the beginning of the diffusion process, and employs VSD loss [50] as a regularization technique to condense a multi-step pre-trained T2I model into a one-step Real-ISR model. However, due to the incorporation of alternating training strategies, OSEDiff may initially tend towards unreliable optimization directions, which may lead to the visual artifacts.
3 Methodology
Figure 2: Pipeline overview. We train one-step Student Model $G_{\theta}$ to map the low-quality image $x_{L}$ into a more realistic one. The noisy latent $\boldsymbol{\hat{z}_{t}}$ and $\boldsymbol{z_{t}}$ sampled by DASM (Details can be found in Fig. 6.) will be fed into both the pre-trained Teacher and LoRA Model to produce the Variational Score loss. Subsequently, two Models’ predictions on $\boldsymbol{z_{t}}$ yield Target Score loss. Their weighted forms, namely TSD ( red flow), in conjunction with pixel-space reconstruction loss ( green flow), are leveraged to update the Student Model. After updating the Student Model, we employ the diffusion loss ( blue flow) to alternately update the LoRA Model.
3.1 Preliminaries
Problem Formulation. ISR problem aims to reconstruct a HQ image ${x}_{H}$ from a LQ input $x_{L}$ by training a parameterized ISR model $G_{\theta}$ on a dataset $\mathcal{D}=\{(x_{L},x_{H})_{i=1}^{N}\}$, where $N$ represents the number of image pairs. It is mathematically formulated as minimizing the following objective:
$\theta^{*}=\arg\min_{\theta}\mathbb{E}_{(x_{L},x_{H})\sim\mathcal{D}}[\mathcal{L}_{Rec}(G_{\theta}(x_{L}),x_{H})$ (1)
$+\lambda\mathcal{L}_{Reg}(q_{\theta}(\hat{x}_{H}),p(x_{H}))]$
Here, $\mathcal{L}_{Rec}$ denotes the reconstruction loss, commonly measured using distance metrics like $L_{2}$ or $LPIPS$ [65]. The regularization term $\mathcal{L}_{Reg}$ enhances the realism and generalization of the ISR model’s outputs. This objective can be understood as aligning the ISR output $\hat{x}_{H}$’s distribution, $q_{\theta}(\hat{x}_{H})$, with the true distribution $p(x_{H})$ by minimizing the KL-divergence [25]:
$\min_{\theta}\mathcal{D}_{\mathrm{KL}}\left(q_{\theta}(\hat{x}_{H})\|p(x_{H})\right)$ (2)
While several studies [46, 62, 45] have employed adversarial loss to optimize this objective, they often encounter issues like mode collapse and training instability. Recent work [52] achieved the state-of-the-art results using Variational Score Distillation (VSD) as the regularization loss to minimizing this objective, which inspires our research.
Variational Score Distillation. Variational Score Distillation (VSD) [50] was initially introduced for text-to-3D generation by distilling a pre-trained text-to-image diffusion model to optimize a single 3D representation [36].
In the VSD framework, a pre-trained diffusion model, represented as $\boldsymbol{\epsilon}_{\psi}$, and its trainable (LoRA [19]) replica $\boldsymbol{\epsilon}_{\phi}$, are used to regularize the generator network $G_{\theta}$. As outlined in ProlificDreamer [50], the gradient with respect to the generator parameters $\boldsymbol{\theta}$ is formulated as follows:
$\begin{split}&\nabla_{\boldsymbol{\theta}}\mathcal{L}_{\mathrm{VSD}}\left(\boldsymbol{\hat{z}},c_{y}\right)\\ &=\mathbb{E}_{t,\epsilon}\left[\omega(t)\left(\boldsymbol{\epsilon}_{\psi}(\boldsymbol{\hat{z}_{t}};t,c_{y})-\boldsymbol{\epsilon}_{\phi}(\boldsymbol{\hat{z}_{t}};t,c_{y})\right)\frac{\partial\boldsymbol{\hat{z}}}{\partial\boldsymbol{\theta}}\right]\end{split}$ (3)
where $\boldsymbol{\hat{z}_{t}}=\alpha_{t}\boldsymbol{\hat{z}}+\sigma_{t}\boldsymbol{\epsilon}$ is the noisy input, $\boldsymbol{\hat{z}}$ is the latent outputted by the generator network $G_{\theta}$, $\boldsymbol{\epsilon}$ is a gaussian noise, and $\alpha_{t},\sigma_{t}$ are the noise-data scaling constants. $c_{y}$ is a text embedding corresponding to a caption that describes the input image, and $w(t)$ is a time-varying weighting function.
3.2 Overview of TSD-SR
As depicted in Fig. 2, our goal is to distill a given pre-trained T2I DMs into a fast one-step Student Model $G_{\theta}$ with Teacher Model $\boldsymbol{\epsilon}_{\psi}$ and trainable LoRA Model $\boldsymbol{\epsilon}_{\phi}$. We denote the latent outputs of the distilled model as $\boldsymbol{\hat{z}_{0}}$, and HQ latent representations as $\boldsymbol{z_{0}}$. Both $\boldsymbol{\hat{z}_{0}}$ and $\boldsymbol{z_{0}}$ will pass through our Distribution-Aware Sampling Module (DASM) to obtain distribution-based samples $\boldsymbol{\hat{z}_{t}}$ and $\boldsymbol{z_{t}}$ (Sec. 3.4). We train $G_{\theta}$ by minimizing the two losses: a reconstruction loss in pixel space to compare the model outputs against the ground truth, and a regularization loss (from Target Score Distillation) to enhances the realism (Sec. 3.3). After updating the Student Model, we update the LoRA Model with diffusion loss. Finally, in Sec. 3.5, we present an overview of all the losses encountered during the training phase.
3.3 Target Score Distillation
Similar to [52], we introduce VSD loss into our work as a regularization loss to enhance the realism and generalization of the $G_{\theta}$ outputs. Upon reviewing VSD Eq. 3, $\boldsymbol{\epsilon}_{\phi}(\boldsymbol{\hat{z}_{t}};t,c_{y})$ represents the gradient direction of the optimization on $G_{\theta}$ noisy outputs $\boldsymbol{\hat{z}_{t}}$, whereas $\boldsymbol{\epsilon}_{\psi}(\boldsymbol{\hat{z}_{t}};t,c_{y})$ corresponds to the gradient direction on noisy HQ latent representations $\boldsymbol{z_{t}}$. The overarching goal of model optimization is to harmonize the predictions from $\boldsymbol{\hat{z}_{t}}$ with those from $\boldsymbol{z_{t}}$. Nonetheless, this strategy encounters hurdles, especially in the early training phase: the quality of synthetic latents $\boldsymbol{\hat{z}_{t}}$ is not high enough for Teacher Model to give a precise prediction. As illustrated in Fig. 3, the Teacher Model finds it challenging to accurately predict the optimization direction for low-quality synthetic latents, with the cosine similarity to the ideal optimization direction only $0.2$ ($0.88$ on HQ latents). This problem can lead to severe visual artifacts, as evident in Fig. 4(a).
Figure 3: A visual comparison of the gradient direction. We set the timestep $t$ to 100 and calculated the cosine similarity between the predicted directions from the Teacher Model and the true direction (towards the HQ data). The prediction direction for $\boldsymbol{z_{t}}$ closely matches the true direction, but not for $\boldsymbol{\hat{z}_{t}}$, suggesting that suboptimal samples may lead to directional deviations.
(a) Naive
(b) MSE
(c) Ours
Figure 4: The visualization of different strategies. (a) The naive method introduces fake textures and fails to recover fine details. (b) MSE leads to over-smoothed generation result, lacking high-frequency information. (c) Our method offers the superior visual effects and fine textures.
A straightforward remedial measure is to employ a mean squared error (MSE) loss to align the synthetic latents with the ideal inputs of the Teacher Model, which are derived from HQ latents. However, as shown in Fig. 4(b), this approach has been observed to lead to over-smoothed results [13]. Our strategy, instead, is to align the predictions made by the Teacher Model on both synthetic and HQ latents, thereby encouraging greater consistency between them. The core idea is that for samples drawn from the same distribution, the real scores predicted by the Teacher Model should be close to each other. We refer to this approach as Target Score Matching (TSM):
$\begin{split}&\nabla_{\theta}\mathcal{L}_{\mathrm{TSM}}(\mathbf{\hat{z}},\mathbf{z},c_{y})\\ &=\mathbb{E}_{t,\epsilon}\left[w(t)({\epsilon}_{\psi}(\mathbf{\hat{z}}_{t};t,c_{y})-{\epsilon}_{\psi}(\mathbf{z}_{t};t,c_{y}))\frac{\partial\mathbf{\hat{z}}}{\partial\theta}\right]\end{split}$ (4)
where the expectation of the gradient is computed across all diffusion timesteps $t\in\{1,\cdots,T\}$ and $\epsilon\sim\mathcal{N}(0,I)$. Equation 4 encapsulates the optimization loss for our Target Score Matching. Upon examining it in conjunction with the Eq. 3, we notice that VSD utilizes the prediction residual between a Teacher and a LoRA Model to drive gradient backpropagation. Conversely, our TSM employs the synthetic and the HQ data to produce the gradients. By blending these two strategies with hyperparameter weighting $\lambda$ and $1-\lambda$, we can construct our ultimate optimization loss, effectively unifying the strengths of both approaches to optimize the training process Eq. 5.
$\begin{split}\nabla_{\theta}\mathcal{L}_{\mathrm{TSD}}(\mathbf{\hat{z}},\mathbf{z},c_{y})=\mathbb{E}_{t,\epsilon}\Big{[}w(t)[{\epsilon}_{\psi}(\mathbf{\hat{z}}_{t};t,c_{y})-\\ {\epsilon}_{\psi}(\mathbf{z}_{t};t,c_{y})+\lambda({\epsilon}_{\psi}(\mathbf{z}_{t};t,c_{y})-{\epsilon}_{\phi}(\mathbf{\hat{z}}_{t};t,c_{y}))]\frac{\partial\mathbf{\hat{z}}}{\partial\theta}\Big{]}\end{split}$ (5)
where $w(t)$ is time-aware weighting function tailored for Real-ISR. Other symbols are in accordance with those previously mentioned. By introducing the prediction of pre-trained diffusion model on HQ latents, we have circumvented the issue of the model falling into the visual artifacts or over-smoothed problem, as illustrated in Fig. 4(c).
3.4 Distribution-Aware Sampling Module
In the VSD-based framework, we need to match the score functions predicted by the Teacher Model with those by the LoRA Model at different timesteps $t\sim\{0,1,\cdots,T\}$. However, for the Real-ISR problem, this matching ability at different $t$ is unbalanced, as illustrated in Fig. 5(a) . This phenomenon could be attributed to our leveraging of the low-frequency (LF) priors from LQ data ,but losing guidance from high-frequency (HF) details. The noisy samples $\hat{z}_{t}$ possesses low-frequency priors (from LQ data). These priors are easily learned by the LoRA Model, leading to more similar predictions (Fig. 5(b)) during LF restoration (Stage 1). However, In the Stage 2, the lack of HF details in $\hat{z}_{t}$ causes the LoRA Model to poorly perceive details, resulting in divergent predictions (Fig. 5(c)). Our goal is to reduce this divergence.
Figure 5: (a) The prediction errors of the VSD loss at different timesteps. The error divergence is more pronounced in early timesteps than later. This phenomenon is observed throughout the optimization process. (b) the visualization of Stage 1 prediction error. (c) the visualization of Stage 2 prediction error.
Existing method, for each iteration, only matches the score function at a single timestep sample $\hat{z}_{t}$, with uniform sampling strategy of time $t$. This leads to slow convergence or even difficulty in optimization during Stage 2, as the gradients of important timesteps are averaged out. To this end, we propose our Distribution-Aware Sampling Module (DASM). This module accumulates optimization gradients for earlier timestep samples in a single iteration, enabling the backpropagation of more gradients focused on detail optimization. As shown in Fig. 6, we first introduce noise into the synthetic latent representation $\boldsymbol{\hat{z}_{t}}=(1-\sigma_{t})\boldsymbol{\hat{z}_{0}}+\sigma_{t}\boldsymbol{\epsilon}$, where $\sigma$ is a weight factor and $\epsilon$ is a gaussian noise. Subsequently, we employ a LoRA Model to perform denoising, yielding the previous time noisy samples by Eq. 6:
$\boldsymbol{\hat{z}_{t-1}}=\boldsymbol{\hat{z}_{t}}+(\sigma_{t-1}-\sigma_{t})\cdot\boldsymbol{\epsilon}_{\phi}(\boldsymbol{\hat{z}_{t}};t,c_{y}),$ (6)
where $\sigma_{t-1}$ and $\sigma_{t}$ are from the flow matching scheduler, and here LoRA Model has fit $\boldsymbol{\hat{z}_{0}}$’s distribution. Similarly, $\boldsymbol{z_{t-1}}$ can be obtained by denoising on the Teacher Model. Ultimately, in a single iteration, the gradients from noise samples along its trajectory can be accumulated to update the Student Model. Since the obtained samples follow the diffusion sampling trajectory and are directed towards early timestep, we effectively optimize the divergence in Stage 2.
Figure 6: Illustration of DASM. Top: The naive method that add noise to samples. Bottom: Our DASM leverages the priors of diffusion models to sample noise that more closely aligns with the true sampling trajectory. The noisy samples can all serve as inputs to the subsequent network to generate gradient backpropagation.
3.5 Training Objective
We summarize all the losses we used in our framework.
Student Model $G_{\theta}$. We train our Student Model with the reconstruction loss $\mathcal{L}_{Rec}$ and the regularization loss $\mathcal{L}_{Reg}$.
For the reconstruction loss, we use $LPIPS$ loss:
$\mathcal{L}_{Rec}\left(G_{\theta}(\boldsymbol{x}_{L}),\boldsymbol{x}_{H}\right)=\mathcal{L}_{LPIPS}\left(G_{\theta}(\boldsymbol{x}_{L}),\boldsymbol{x}_{H}\right).$ (7)
For the regularization loss, we use our TSD loss Eq. 5. Therefore, the overall training objective for the Student Model $G_{\theta}$ is:
$\mathcal{L}_{Stu}=\mathcal{L}_{Rec}+\gamma\mathcal{L}_{Reg},$ (8)
where $\gamma$ is a weighting scalar.
LoRA Model $\boldsymbol{\epsilon}_{\phi}$. As stipulated by VSD, the fine-tuned replica $\epsilon_{\phi}$ must be trainable, with its training objective being:
$\mathcal{L}_{Diff}(\boldsymbol{{\hat{z}}},c_{y})=\mathbb{E}_{t,\boldsymbol{\epsilon}}[\left\|\boldsymbol{\epsilon}_{\phi}(\boldsymbol{\hat{z}_{t}};t,c_{y})-\boldsymbol{\epsilon}^{\prime}\right\|^{2}],$ (9)
where $\boldsymbol{\epsilon}^{\prime}$ serves as the training target for the denoising network, representing Gaussian noise in the context of DDPM, and a gradient towards HQ for flow matching.
4 Experiments
4.1 Experimental Settings
Training Datasets. We adopt DIV2K [1], Flickr2K [42], LSDIR [27] and first 10K face images from FFHQ [20] for training. We use the same degradation pipeline as Real-ESRGAN [46] to synthesize LR-HR pairs.
Test Datasets. We evaluated our model on synthetic DIV2K-Val [1] dataset and two real-world datasets, RealSR [4] and DRealSR [51]. The real-world datasets comprised 128x128 LQ and 512x512 HQ image pairs. The synthetic dataset included 3,000 image pairs, which are created by downgrading the 512x512 HQ images croped from DIV2K-Val to 128x128 LQ images using the Real-ESRGAN [46] degradation pipeline.
Evaluation Metrics. For evaluating our method, we apply both full-reference and no-reference metrics. Full-reference metrics include PSNR and SSIM [49] (calculated on the Y channel in YCbCr space) for fidelity, LPIPS [65] and DISTS [7] for perceptual quality, and FID [14] for distribution comparison. No-reference metrics include NIQE [63], MANIQA [55], MUSIQ [22] and CLIPIQA [43].
Compared Methods. We categorize the test models into two groups: single-step and multi-step inference. The single-step inference diffusion models include SinSR [48], AddSR [54], and OSEDiff [52]. The multi-step inference diffusion models comprise StableSR [44], ResShift [61], PASD [56], DiffBIR [29], SeeSR [53], SUPIR [60], and AddSR [54]. Specifically, for AddSR, we have conducted comparisons between its single-step and four-step models. GAN-based Real-ISR methods [62, 5, 28, 46] are detailed in the supplementary material.
Implementation Details. All models mentioned are initialized from the Teacher Model (SD3 [10] in our work). We only train vae encoder and denoising network in the Student Model, frozen the vae decoder in order to preserve vae’s prior [24]. We utilize the default prompt for the Student Model, and prompts extracted from HQ for the Teacher and LoRA models when training. We use the AdamW optimizer [30] with a learning rate of 5e-6 for the Student Model and 1e-6 for the LoRA Model, and set rank of LoRA to 64 for both two models. The training process took roughly 48h, utilizing 8 NVIDIA V100 GPUs with a batch size of 16.
Table 1: Quantitative comparison with the state-of-the-art one-step methods across both synthetic and real-world benchmarks. The number of diffusion inference steps is indicated by ‘s’. The best results of each metric are highlighted in red.
Datasets Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ NIQE $\downarrow$ MUSIQ $\uparrow$ MANIQA $\uparrow$ CLIPIQA $\uparrow$
OSEDiff-1s 27.35 0.7610 0.3177 0.2365 141.93 7.3053 63.56 0.5763 0.7053
AddSR-1s 27.77 0.7722 0.3196 0.2242 150.18 6.9321 60.85 0.5490 0.6188
SinSR-1s 28.38 0.7497 0.3669 0.2484 172.72 6.9606 55.03 0.4904 0.6412
DRealSR Ours-1s 26.01 0.7174 0.3092 0.2216 130.83 5.5044 66.66 0.5850 0.7345
OSEDiff-1s 23.94 0.6736 0.3172 0.2363 125.93 6.3822 67.52 0.6187 0.7001
AddSR-1s 24.79 0.7077 0.3091 0.2191 132.05 5.5440 66.18 0.6098 0.5722
SinSR-1s 26.27 0.7351 0.3217 0.2341 137.59 6.2964 60.76 0.5418 0.6163
RealSR Ours-1s 23.27 0.6874 0.2807 0.2185 113.31 5.0679 71.28 0.6346 0.7235
OSEDiff-1s 23.72 0.6109 0.3058 0.2138 26.34 5.3903 65.27 0.5838 0.6558
AddSR-1s 23.26 0.5902 0.3623 0.2123 29.68 4.7610 63.39 0.5657 0.5734
SinSR-1s 24.41 0.6018 0.3262 0.2068 35.55 5.9981 62.95 0.5430 0.6501
DIV2K-Val Ours-1s 22.12 0.5595 0.2704 0.1893 23.55 4.3730 71.38 0.6136 0.7273
4.2 Comparison with Existing Methods
Quantitative Comparisons. Tab. 1 show the quantitative comparison of our method with single-step diffusion models on three datasets. Ours achieves the best results on all metrics except PSNR and SSIM. SinSR and AddSR, as distilled versions of previous multi-step super-resolution methods, reduce inference steps but experience a corresponding decrease in performance metrics. OSEDiff introduces the VSD loss from 3D generation tasks into Real-ISR without fully addressing the substantial differences between the two domains. Therefore, the no-reference metrics for image quality are not very satisfied. In contrast, our proposed TSD-SR, tailored for Real-ISR, outperform other single-step models in terms of the vast majority of key metrics.
Tab. 2 show the quantitative comparison with multi-step models. We can draw the following conclusions: (1) TSD-SR demonstrates significant advantages over competing methods in terms of LPIPS, FID, and NIQE metrics. Additionally, it achieves performance that surpasses most multi-step models on DISTS, MUSIQ, and CLIPIQA. (2) DiffBIR, SeeSR, PASD, and AddSR exhibit better performance on the MANIQA metric, which may be attributed to the fact that multi-step models have more denoising iterations to produce rich details. (3) ResShift stands out with the highest PSNR and SSIM scores, while StableSR also shows notable performance regarding DISTS and FID metrics. However, both two models underperform in terms of no-reference metrics.
Finally, we explain the lower PSNR and SSIM metrics in our experiments. Several works [54, 60] have found that these reconstruction metrics are not well-suited for evaluating Real-ISR tasks. When a model recovers better details, it leads to lower reconstruction metrics, indicating a trade-off [3, 67, 32]. Refer to the supplementary material for detailed visual comparisons and analysis.
Table 2: Quantitative comparison with state-of-the-art multi-step methods across both synthetic and real-world benchmarks. The number of diffusion inference steps is indicated by ‘s’. The best and second best results of each metric are highlighted in red and blue, respectively.
Datasets Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ NIQE $\downarrow$ MUSIQ $\uparrow$ MANIQA $\uparrow$ CLIPIQA $\uparrow$
StableSR-200s 28.04 0.7454 0.3279 0.2272 144.15 6.5999 58.53 0.5603 0.6250
DiffBIR-50s 25.93 0.6525 0.4518 0.2761 177.04 6.2324 65.66 0.6296 0.6860
SeeSR-50s 28.14 0.7712 0.3141 0.2297 146.95 6.4632 64.74 0.6022 0.6893
SUPIR-50s 25.09 0.6460 0.4243 0.2795 169.48 7.3918 58.79 0.5471 0.6749
PASD-20s 27.79 0.7495 0.3579 0.2524 171.03 6.7661 63.23 0.5919 0.6242
ResShift-15s 28.69 0.7874 0.3525 0.2541 176.77 7.8762 52.40 0.4756 0.5413
AddSR-4s 26.72 0.7124 0.3982 0.2711 164.12 7.6689 66.33 0.6257 0.7226
DRealSR Ours-1s 26.01 0.7174 0.3092 0.2216 130.83 5.5044 66.66 0.5850 0.7345
StableSR-200s 24.62 0.7041 0.3070 0.2156 128.54 5.7817 65.48 0.6223 0.6198
DiffBIR-50s 24.24 0.6650 0.3469 0.2300 134.56 5.4932 68.35 0.6544 0.6961
SeeSR-50s 25.21 0.7216 0.3003 0.2218 125.10 5.3978 69.69 0.6443 0.6671
SUPIR-50s 23.65 0.6620 0.3541 0.2488 130.38 6.1099 62.09 0.5780 0.6707
PASD-20s 25.68 0.7273 0.3144 0.2304 134.18 5.7616 68.33 0.6323 0.5783
ResShift-15s 26.39 0.7567 0.3158 0.2432 149.59 6.8746 60.22 0.5419 0.5496
AddSR-4s 23.33 0.6400 0.3925 0.2626 154.22 5.8959 71.49 0.6826 0.7225
RealSR Ours-1s 23.27 0.6874 0.2807 0.2185 113.31 5.0679 71.28 0.6346 0.7235
StableSR-200s 23.27 0.5722 0.3111 0.2046 24.95 4.7737 65.78 0.6164 0.6753
DiffBIR-50s 23.13 0.5717 0.3469 0.2108 33.93 4.6056 68.54 0.6360 0.7125
SeeSR-50s 23.73 0.6057 0.3198 0.1953 25.81 4.8322 68.49 0.6198 0.6899
SUPIR-50s 22.13 0.5279 0.3919 0.2312 31.40 5.6767 63.86 0.5903 0.7146
PASD-20s 24.00 0.6041 0.3779 0.2305 39.12 4.8587 67.36 0.6121 0.6327
ResShift-15s 24.71 0.6234 0.3473 0.2253 42.01 6.3615 60.63 0.5283 0.5962
AddSR-4s 22.16 0.5280 0.4053 0.2360 35.41 5.2584 70.99 0.6596 0.7593
DIV2K-Val Ours-1s 22.12 0.5595 0.2704 0.1893 23.55 4.3730 71.38 0.6136 0.7273
Qualitative Comparisons. Fig. 7 present visual comparisons of different Real-ISR methods. As shown in multi-step methods’ results, SeeSR uses degradation-aware semantic cues to leverage image generation priors, but it sometimes produces over-smooth textures. SUPIR demonstrates a notably robust generative capacity, however, the overproduction of extraneous details will lead to results that lack naturalism in image restoration (e.g. Adding excessive wrinkles at the corners of a young girl’s eyes). Under more realistic degradation conditions, PASD finds it challenging to restore the corresponding content, revealing a deficiency in the model’s robustness capabilities. Among single-step methods, SinSR introduces artifacts, possibly because its distilled pre-trained model is trained from scratch, lacking sufficient real-world priors, which results in unsatisfactory image restoration capabilities. AddSR produces over-smoothed results when using its 1-step model. OSEDiff offers improved restoration effects compared to SinSR and AddSR, yet it may fall short in terms of authenticity and naturalness when it comes to detail recovery. In contrast, our method effectively generates rich textures and realistic details with enhanced sharpness and contrast. Additional visual comparisons and results are provided in the supplementary material.
Figure 7: Visual comparisons of different Real-ISR methods. Please zoom in for a better view.
Complexity Comparisons
Table 3: Comparison of computational complexity across different DMs-based methods. The performance of each method is measured on an A100 GPU, using input images sized at 512 × 512 pixels. We have disregarded the loading time for model weights and data.
StableSR DiffBIR SeeSR SUPIR PASD ResShift AddSR AddSR SinSR OSEDiff Ours
Inference Step 200 50 50 50 20 15 4 1 1 1 1
Inference Time 12.4151 7.9637 5.8167 16.8704 4.8441 0.7546 1.0199 0.5043 0.1424 0.1650 0.1362
We assess the complexity of the state-of-the-art DM-based Real-ISR models as detailed in Tab. 3, focusing on inference time. The performance of each method is benchmarked on an A100 GPU with input images sized at 512 × 512 pixels. We have disregarded the loading time for model weights and data, and the main computation time includes: (1) text extraction time when a text extractor is used; (2) text encoder computation time if needed; (3) VAE encoding and decoding time; (4) denoising network execution time. It is evident that TSD-SR has a substantial advantage in inference time compared with multi-step models. Specifically, TSD-SR is more than 120 times faster than SUPIR, 90 times faster than StableSR, about 50 times faster than DiffBIR, more than 40 times faster than SeeSR, more than 35 times faster than PASD, and 4 times faster than ResShift. When compared with existing one-step models, our model boasts the fastest inference times. This is because we directly denoise from LQ data and use a fixed prompt.
4.3 User Study
We conduct a user study comparing our method with three other diffusion-based one-step super-resolution methods. To make the evaluation more comprehensive, we selected images from five categories—human faces, buildings, animals, vegetation and characters. A total of 50 participants were engaged in the voting process. We guided participants to evaluate the best restoration results based on the similarity to HQ image, structural similarity to the LQ image, and realism of textures and details. The results shown in the Fig. 8 indicate that our method receives a 69.2% approval rate from users. We scored 57.60% in Animals, 70.00% in Buildings, 68.80% in Human Faces, 65.20% in Vegetation, and 84.40% in Characters, surpassing other models.
Figure 8: The results of our user study. Left: Category-based user preference radar chart. Our model has won the most user favor across all categories. Right: User preference pie chart. Our approach has garnered a 69.2% user satisfaction rating.
Table 4: Ablation study of Target Score Matching loss and Distribution-Aware Sampling Module.
Datasets Method NIQE $\downarrow$ MUSIQ $\uparrow$ MANIQA $\uparrow$ CLIPIQA $\uparrow$
DRealSR w/o TSM 5.6268 63.90 0.5749 0.6958
w/o DASM 6.0368 63.56 0.5812 0.7123
Full 5.5044 66.66 0.5850 0.7345
RealSR w/o TSM 5.2728 68.56 0.6338 0.6987
w/o DASM 5.5638 69.05 0.6273 0.7031
Full 5.0679 71.28 0.6346 0.7235
4.4 Ablation Study
To validate the effectiveness of TSM loss and DASM, we conducted ablation studies by removing them separately in our experiments. We selected no-reference metrics for comparison, as these are critical for assessing image quality. The results are presented in Tab. 4. We draw the following conclusions: (1) The absence of TSM loss and DASM affects the no-reference metrics NIQE, MUSIQ, MANIQA, and CLIPIQA. (2) The lack of TSM leads to a significant decrease in MUSIQ and CLIPIQA metrics, possibly due to the unreliable directions in VSD causing unrealistic generations. (3) The absence of DASM results in a decline in NIQE, MUSIQ and CLIPIQA, possibly due to the suboptimal optimization of details.
5 Conclusion and Limitation
We propose TSD-SR, an effective and one-step model based on diffusion prior for Real-ISR. TSD-SR utilizes the TSD to enhance the realism of images generated by the distillation model. And it leverages DASM to sample distribution-based samples and accumulate their gradients to enhance the recovery of details. Our experiments have demonstrated that TSD-SR outperforms existing one-step Real-ISR models in both performance and inference speed.
limitations. Although our model boasts excellent inference speed and restoration performance, it still has a large number of model parameters compared to the past GAN/non-diffusion approaches. In the future, we plan to employ pruning or quantization methods to compress the model parameters, striving for a lightweight and efficient Real-ISR model.
Supplementary Material
Appendix A Comparison with GAN-based Methods
We compared our method with GAN-based approaches in Tab. 5. While the GAN methods show advantages in full-reference metrics such as PSNR and SSIM, our model outperforms GAN-based methods on all no-reference metrics. Some research has found the limitations of full-reference methods in the field of image super-resolution [60, 54]. Full-reference metrics often assign low scores to visually appealing results because these results produce more details, some of which may not align with the ground truth. However, no-reference metrics assess image quality based on the individual image, without the need to forcibly align with the ground truth. Therefore, in more complex and realistic degradation scenarios, no-reference metrics may be more suitable for evaluating the results of image super-resolution. In Appendix C, we discuss the comparison between full-reference metrics and human preferences, and in the Fig. 9, we present a visual comparison with GAN-based methods. From the visualization, it can be observed that our model achieves better results in terms of texture details compared to GAN methods.
Table 5: Quantitative comparison with GAN-based methods on both synthetic and real-world benchmarks. The best results of each metric are highlighted in red.
Datasets Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ NIQE $\downarrow$ MUSIQ $\uparrow$ MANIQA $\uparrow$ CLIPIQA $\uparrow$
BSRGAN 28.70 0.8028 0.2858 0.2143 155.61 6.5408 57.15 0.4847 0.5091
Real-ESRGAN 28.61 0.8051 0.2818 0.2088 147.66 6.7001 54.27 0.4888 0.4521
LDL 28.20 0.8124 0.2791 0.2127 155.51 7.1448 53.94 0.4894 0.4476
FeMASR 26.87 0.7569 0.3156 0.2238 157.72 5.9067 53.70 0.4413 0.5633
DRealSR Ours 26.01 0.7174 0.3092 0.2216 130.83 5.5044 66.66 0.5850 0.7345
BSRGAN 26.38 0.7651 0.2656 0.2123 141.24 5.6431 63.28 0.5425 0.5114
Real-ESRGAN 26.65 0.7603 0.2726 0.2065 136.29 5.8471 60.45 0.5507 0.4518
LDL 25.28 0.7565 0.2750 0.2119 142.74 5.9880 60.92 0.5494 0.4559
FeMASR 25.06 0.7356 0.2936 0.2285 141.01 5.7696 59.05 0.4872 0.5405
RealSR Ours 23.27 0.6874 0.2807 0.2185 113.31 5.0679 71.28 0.6346 0.7235
BSRGAN 24.58 0.6269 0.3502 0.2280 49.55 4.7501 61.68 0.4979 0.5386
Real-ESRGAN 24.02 0.6387 0.3150 0.2123 38.87 4.8271 60.38 0.5401 0.5251
LDL 23.83 0.6344 0.3256 0.2227 42.28 4.8555 60.04 0.5328 0.5180
FeMASR 22.45 0.5858 0.3370 0.2205 41.97 4.8679 57.94 0.4787 0.5769
DIV2K-Val Ours 22.12 0.5595 0.2704 0.1893 23.55 4.3730 71.38 0.6136 0.7273
Figure 9: Qualitative comparisons between TSD-SR and GAN-based Real-ISR methods. Please zoom in for a better view.
Appendix B More Visual Comparisons
In Figs. 10, 11 and 12, we provide more visual caparisons with other diffusion-based methods. Numerous examples demonstrate the robust restoration capabilities of TSD-SR and the high quality of the restored images.
Figure 10: Qualitative comparisons between TSD-SR and different diffusion-based methods. Our method can effectively restore the texture and details of the corresponding object under challenging degradation conditions. Please zoom in for a better view.
Figure 11: Qualitative comparisons between TSD-SR and different diffusion-based methods. Our method can effectively restore the texture and details of the corresponding object under challenging degradation conditions. Please zoom in for a better view.
Figure 12: Qualitative comparisons between TSD-SR and different diffusion-based methods. Our method can effectively restore the texture and details of the corresponding object under challenging degradation conditions. Please zoom in for a better view.
Appendix C Comparisons of Full-reference Metrics and Human Preference
We present additional comparative experiments to demonstrate the limitations of full-reference metrics in Figure 13. It can be observed that GAN-based methods with advantages in PSNR and SSIM often result in blurred details and textures. In contrast, our results offer better visual effects, yet they do not hold an advantage in terms of PSNR and SSIM metrics. This phenomenon has also been discussed in other research works [3, 67, 32], with the majority of studies suggesting that for models with strong generation and restoration capabilities, these full-reference metrics do not adequately reflect image quality. We anticipate the development of better full-reference metrics in the future to assess advanced Real-ISR methods.
Figure 13: Comparisons between full-reference metric assessments and human visual preference. Despite scoring lower on full-reference metrics, TSD-SR generates images that align with human preference.
Appendix D Theory of Target Score Matching
The core idea of Target Score Matching (TSM) is that for samples drawn from the same distribution, the real scores predicted by the Teacher Model should be close to each other. Thus we minimize the MSE loss between the Teacher Model’s predictions of $\mathbf{\hat{z}}_{t}$ and $\mathbf{z}_{t}$ by
$\begin{split}&\mathcal{L}_{\mathrm{MSE}}(\mathbf{\hat{z}},\mathbf{z},c_{y})\\ &=\mathbb{E}_{t,\epsilon}\left[w(t)\|\epsilon_{\psi}(\mathbf{\hat{z}}_{t};t,c_{y})-\epsilon_{\psi}(\mathbf{z}_{t};t,c_{y})\|_{2}^{2}\right]\end{split}$ (10)
where the expectation of the gradient is computed across all diffusion timesteps $t\in\{1,\cdots,T\}$ and $\epsilon\sim\mathcal{N}(0,I)$.
To understand the difficulties of this approach, consider the gradient of
$\begin{split}&\nabla_{\theta}\mathcal{L}_{\mathrm{MSE}}(\mathbf{\hat{z}},\mathbf{z},c_{y})=\mathbb{E}_{t,\epsilon}\Big{[}w(t)\cdot\underbrace{\frac{\partial{\epsilon}_{\psi}(\mathbf{\hat{z}}_{t};t,c_{y})}{\partial\mathbf{\hat{z}}_{t}}}_{\text{Diffusion Jacobian}}\\ &\underbrace{({\epsilon}_{\psi}(\mathbf{\hat{z}}_{t};t,c_{y})-{\epsilon}_{\psi}(\mathbf{z}_{t};t,c_{y}))}_{\text{Prediction Residual}}\underbrace{\frac{\partial\mathbf{\hat{z}}}{\partial\theta}}_{\text{Generator Jacobian}}\Big{]}\end{split}$ (11)
where we absorb $\frac{\partial\mathbf{\mathbf{\hat{z}}}_{t}}{\partial\mathbf{\mathbf{\hat{z}}}}$ and the other constant into $w(t)$. The computation of the Diffusion Jacobian term is computationally demanding, as it necessitates backpropagation through the Teacher Model. DreamFusion [36] found that this term struggles with small noise levels due to its training to approximate the scaled Hessian of marginal density. This work also demonstrated that omitting the Diffusion Jacobian term leads to an effective gradient for optimizing. Similar to their approach, we update Eq. 11 by omitting Diffusion Jacobian:
$\begin{split}&\nabla_{\theta}\mathcal{L}_{\mathrm{TSM}}(\mathbf{\hat{z}},\mathbf{z},c_{y})=\\ &\mathbb{E}_{t,\epsilon}\Big{[}w(t)\underbrace{({\epsilon}_{\psi}(\mathbf{\hat{z}}_{t};c_{y},t)-{\epsilon}_{\psi}(\mathbf{z}_{t};c_{y},t))}_{\text{Prediction Residual}}\underbrace{\frac{\partial\mathbf{\hat{z}}}{\partial\theta}}_{\text{Generator Jacobian}}\Big{]}\end{split}$ (12)
The effectiveness of the method can be proven by starting from the KL divergence. We can use a Sticking-the-Landing style gradient by thinking of ${\epsilon}_{\psi}(\mathbf{z}_{t};c_{y},t)$ as a control variate for $\hat{\epsilon}$ For detailed proof, refer to Appendix 4 of DreamFusion [36]. It demonstrates that the gradient of this loss yields the same updates as optimizing the training loss $\mathcal{L}_{\mathrm{MSE}}$ Eq. 10, excluding the Diffusion Jacobian term.
Compared with the VSD loss, we find that the term “Prediction Residual” has changed, and the two losses are similar in the gradient update mode. Specifically, we find that VSD employs identical inputs for both the Teacher and LoRA models to compute the gradient, while here TSM uses high-quality and suboptimal inputs for the Teacher Model. The losses are related to each other through $\epsilon_{\phi}(\mathbf{\hat{z}}_{t};t,c_{y})$.
Appendix E Algorithm
Algorithm 1 details our TSD-SR training procedure. We use classifier-free guidance (cfg) for the Teacher Model and the LoRA Model. The cfg value is set to 7.5.
Input: $\mathcal{D}$ = $\{x_{L},x_{H},c_{y}\}$ , pre-trained Teacher Diffusion Model including VAE encoder $E_{\psi}$, denoising network $\epsilon_{\psi}$ and VAE decoder $D_{\psi}$, the number of iterations $N$ and step size $s$ of DASM.
Output: Trained one-step Student Model $G_{\theta}$.
1 Initialize Student Model $G_{\theta}$, including $E_{\theta}\leftarrow E_{\psi}$ with trainable LoRA, $\epsilon_{\theta}\leftarrow\epsilon_{\psi}$ with trainable LoRA, $D_{\theta}\leftarrow D_{\psi}$.
2Initialize LoRA diffusion network $\epsilon_{\phi}\leftarrow\epsilon_{\psi}$ with trainable LoRA.
3while train do
4       Sample $(x_{L},x_{H},c_{y})\sim\mathcal{D}$
      /* Network forward */
5
6      $\boldsymbol{\hat{z}}\leftarrow\epsilon_{\theta}(E_{\theta}(x_{L}))$, $\boldsymbol{z}\leftarrow E_{\psi}(x_{H})$
7      $\hat{x}_{H}\leftarrow D_{\psi}(\boldsymbol{\hat{z}})$
      /* Compute reconstruction loss */
8
9      $\mathcal{L}_{Rec}\leftarrow LPIPS(\hat{x}_{H},x_{H})$
      /* Compute regularization loss */
10
11      Sample $\epsilon$ from $\mathcal{N}(0,I)$, $t$ from $\{50,\cdots,950\}$
12      $\sigma_{t}\leftarrow$ FlowMatchingScheduler($t$)
13      $\boldsymbol{\hat{z}}_{t}\leftarrow\sigma_{t}\epsilon+(1-\sigma_{t})\boldsymbol{\hat{z}}$, $\boldsymbol{z}_{t}\leftarrow\sigma_{t}\epsilon+(1-\sigma_{t})\boldsymbol{z}$
14      $\mathcal{L}_{Reg}\leftarrow\mathcal{L}_{TSD}(\boldsymbol{\hat{z}}_{t},\boldsymbol{z}_{t},c_{y})$ // Eq. 5
15      for $i\leftarrow 1$ to $N$ do
16             $cur\leftarrow t-i\cdot s$
17            $pre\leftarrow t-i\cdot s+s$
18            $\sigma_{cur}\leftarrow$ FlowMatchingScheduler($cur$)
19            $\sigma_{pre}\leftarrow$ FlowMatchingScheduler($pre$)
20            $\boldsymbol{\hat{z}}_{cur}\leftarrow\boldsymbol{\hat{z}}_{pre}+(\sigma_{cur}-\sigma_{pre})\cdot\epsilon_{\phi}(\boldsymbol{\hat{z}}_{pre};pre,c_{y})$ // Eq. 6
21            $\boldsymbol{z}_{cur}\leftarrow\boldsymbol{z}_{pre}+(\sigma_{cur}-\sigma_{pre})\cdot\epsilon_{\psi}(\boldsymbol{z}_{pre};pre,c_{y})$
22            $\mathcal{L}_{Reg}$ += $weight\cdot\mathcal{L}_{TSD}(\boldsymbol{\hat{z}}_{cur},\boldsymbol{z}_{cur},c_{y})$
23       end for
24
25      $\mathcal{L}_{G}\leftarrow\mathcal{L}_{Rec}+\gamma\mathcal{L}_{Reg}$
26      Update $\theta$ with $\mathcal{L}_{G}$
      /* Compute diffusion loss for LoRA Model */
27
28      Sample $\epsilon$ from $\mathcal{N}(0,I)$, $t$ from $\{50,\cdots,950\}$
29      $\sigma_{t}\leftarrow$ FlowMatchingScheduler($t$)
30      $\boldsymbol{\hat{z}}_{t}\leftarrow\sigma_{t}\epsilon+(1-\sigma_{t})stopgrad(\boldsymbol{\hat{z}})$
31      $\mathcal{L}_{Lora}\leftarrow\mathcal{L}_{Diff}(\boldsymbol{\hat{z}}_{t},c_{y})$ // Eq. 9
32      Update $\phi$ with $\mathcal{L}_{Lora}$
33 end while
Algorithm 1 TSD-SR Training Procedure

### v3

@@
 3.5 Training Objective
-We summarize all the losses we used in our framework.
+We summarize all the losses that we used in our framework.
 Student Model $G_{\theta}$. We train our Student Model with the reconstruction loss $\mathcal{L}_{Rec}$ and the regularization loss $\mathcal{L}_{Reg}$.
-For the reconstruction loss, we use $LPIPS$ loss:
-$\mathcal{L}_{Rec}\left(G_{\theta}(\boldsymbol{x}_{L}),\boldsymbol{x}_{H}\right)=\mathcal{L}_{LPIPS}\left(G_{\theta}(\boldsymbol{x}_{L}),\boldsymbol{x}_{H}\right).$ (7)
+For the reconstruction loss, we use $LPIPS$ loss and $MSE$ loss in the latent space:
+$\begin{split}&\mathcal{L}_{Rec}\left(G_{\theta}(\boldsymbol{x}_{L}),\boldsymbol{x}_{H}\right)\\ &=\gamma_{1}\mathcal{L}_{LPIPS}\left(G_{\theta}(\boldsymbol{x}_{L}),\boldsymbol{x}_{H}\right)+\mathcal{L}_{MSE}\left(\mathbf{z}_{t},\mathbf{\hat{z}}_{t}\right).\end{split}$ (7)
 For the regularization loss, we use our TSD loss Eq. 5. Therefore, the overall training objective for the Student Model $G_{\theta}$ is:
-$\mathcal{L}_{Stu}=\mathcal{L}_{Rec}+\gamma\mathcal{L}_{Reg},$ (8)
-where $\gamma$ is a weighting scalar.
+$\mathcal{L}_{Stu}=\mathcal{L}_{Rec}+\gamma_{2}\mathcal{L}_{Reg},$ (8)
+where $\gamma_{1}$ and $\gamma_{2}$ are weighting factors.
+We initialize both $\gamma_{1}$ and $\gamma_{2}$ to 1 at the beginning of training.
+As optimization progresses, we ramp up $\gamma_{1}$ from 1 to 2 while maintaining $\gamma_{2}$ at its initial value.
 LoRA Model $\boldsymbol{\epsilon}_{\phi}$. As stipulated by VSD, the fine-tuned replica $\epsilon_{\phi}$ must be trainable, with its training objective being:
@@
 All models mentioned are initialized from the Teacher Model (SD3 [10] in our work).
-We only train vae encoder and denoising network in the Student Model, frozen the vae decoder in order to preserve vae’s prior [24].
-We utilize the default prompt for the Student Model, and prompts extracted from HQ for the Teacher and LoRA models when training.
-We use the AdamW optimizer [30] with a learning rate of 5e-6 for the Student Model and 1e-6 for the LoRA Model, and set rank of LoRA to 64 for both two models.
-The training process took roughly 48h, utilizing 8 NVIDIA V100 GPUs with a batch size of 16.
+Similar to OSEDiff [53], we only train the vae encoder and denoising network in the Student Model, frozen the vae decoder in order to preserve vae’s prior [24].
+We utilize the default prompt for the Student Model, and prompts are extracted from HQ for the Teacher and LoRA models when training.
+We use AdamW optimizer [30] with a learning rate of 5e-6 for the Student Model and 1e-6 for the LoRA Model, and set the rank of LoRA to 64 for both two models.
+During the initial training phase, we incorporate MSE loss in the latent space and exclude DASM to enhance model stability and reduce training time.
+In later stages, we remove this MSE loss as it leads to over-smoothed results, and we employ DASM to improve image restoration.
+The training process took roughly 96h, utilizing 8 NVIDIA V100 GPUs with a batch size of 16.
 Table 1: Quantitative comparison with the state-of-the-art one-step methods across both synthetic and real-world benchmarks.
@@
 Datasets Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ NIQE $\downarrow$ MUSIQ $\uparrow$ MANIQA $\uparrow$ CLIPIQA $\uparrow$
-OSEDiff-1s 27.35 0.7610 0.3177 0.2365 141.93 7.3053 63.56 0.5763 0.7053
+OSEDiff-1s 27.92 0.7836 0.2968 0.2162 135.51 6.4471 64.69 0.5898 0.6958
 AddSR-1s 27.77 0.7722 0.3196 0.2242 150.18 6.9321 60.85 0.5490 0.6188
 SinSR-1s 28.38 0.7497 0.3669 0.2484 172.72 6.9606 55.03 0.4904 0.6412
-DRealSR Ours-1s 26.01 0.7174 0.3092 0.2216 130.83 5.5044 66.66 0.5850 0.7345
-OSEDiff-1s 23.94 0.6736 0.3172 0.2363 125.93 6.3822 67.52 0.6187 0.7001
+DRealSR Ours-1s 27.77 0.7559 0.2967 0.2136 134.98 5.9131 66.62 0.5874 0.7344
+OSEDiff-1s 25.15 0.7341 0.2920 0.2128 123.48 5.6471 69.10 0.6326 0.6687
 AddSR-1s 24.79 0.7077 0.3091 0.2191 132.05 5.5440 66.18 0.6098 0.5722
 SinSR-1s 26.27 0.7351 0.3217 0.2341 137.59 6.2964 60.76 0.5418 0.6163
-RealSR Ours-1s 23.27 0.6874 0.2807 0.2185 113.31 5.0679 71.28 0.6346 0.7235
-OSEDiff-1s 23.72 0.6109 0.3058 0.2138 26.34 5.3903 65.27 0.5838 0.6558
+RealSR Ours-1s 24.81 0.7172 0.2743 0.2104 114.45 5.1298 71.19 0.6347 0.7160
+OSEDiff-1s 23.72 0.6109 0.2942 0.1975 26.34 4.7089 67.96 0.6131 0.6681
 AddSR-1s 23.26 0.5902 0.3623 0.2123 29.68 4.7610 63.39 0.5657 0.5734
 SinSR-1s 24.41 0.6018 0.3262 0.2068 35.55 5.9981 62.95 0.5430 0.6501
-DIV2K-Val Ours-1s 22.12 0.5595 0.2704 0.1893 23.55 4.3730 71.38 0.6136 0.7273
+DIV2K-Val Ours-1s 23.02 0.5808 0.2673 0.1821 29.16 4.3244 71.69 0.6192 0.7416
 4.2 Comparison with Existing Methods
@@
 1 show the quantitative comparison of our method with single-step diffusion models on three datasets.
-Ours achieves the best results on all metrics except PSNR and SSIM.
+Ours achieves the best results on most evaluation metrics.
 SinSR and AddSR, as distilled versions of previous multi-step super-resolution methods, reduce inference steps but experience a corresponding decrease in performance metrics.
@@
 Therefore, the no-reference metrics for image quality are not very satisfied.
-In contrast, our proposed TSD-SR, tailored for Real-ISR, outperform other single-step models in terms of the vast majority of key metrics.
+In contrast, our proposed TSD-SR, tailored for Real-ISR, outperforms other single-step models in terms of the vast majority of key metrics.
 Tab.
 2 show the quantitative comparison with multi-step models.
-We can draw the following conclusions: (1) TSD-SR demonstrates significant advantages over competing methods in terms of LPIPS, FID, and NIQE metrics.
-Additionally, it achieves performance that surpasses most multi-step models on DISTS, MUSIQ, and CLIPIQA.
+We can draw the following conclusions: (1) TSD-SR demonstrates significant advantages over competing methods in terms of LPIPS, DISTS, and NIQE metrics.
+Additionally, it achieves performance that surpasses most multi-step models on FID, MUSIQ, and CLIPIQA.
 (2) DiffBIR, SeeSR, PASD, and AddSR exhibit better performance on the MANIQA metric, which may be attributed to the fact that multi-step models have more denoising iterations to produce rich details.
@@
 When a model recovers better details, it leads to lower reconstruction metrics, indicating a trade-off [3, 68, 32].
-Refer to the supplementary material for detailed visual comparisons and analysis.
+This phenomenon has also been widely discussed in other related research work [3, 68, 32, 61, 66, 45, 55].
+LPIPS [66] is proposed to overcome the limitation that PSNR and SSIM fail to align with human judgments in spatial ambiguities situation.
+Other DMs-based SR researchers [61, 45] argue that DMs introduce superior pre-trained priors, enabling the restoration of information that traditional methods (from scratch) cannot achieve.
+However, such capability often leads to a decline in pixel-level metrics, as they prioritize distribution modeling and sampling from learned distributions over strict pixel fidelity.
+We anticipate the development of better full-reference metrics in the future to assess advanced Real-ISR methods.
+Refer to the supplementary material for detailed visual comparisons.
 Table 2: Quantitative comparison with state-of-the-art multi-step methods across both synthetic and real-world benchmarks.
@@
 AddSR-4s 26.72 0.7124 0.3982 0.2711 164.12 7.6689 66.33 0.6257 0.7226
-DRealSR Ours-1s 26.01 0.7174 0.3092 0.2216 130.83 5.5044 66.66 0.5850 0.7345
+DRealSR Ours-1s 27.77 0.7559 0.2967 0.2136 134.98 5.9131 66.62 0.5874 0.7344
 StableSR-200s 24.62 0.7041 0.3070 0.2156 128.54 5.7817 65.48 0.6223 0.6198
@@
 AddSR-4s 23.33 0.6400 0.3925 0.2626 154.22 5.8959 71.49 0.6826 0.7225
-RealSR Ours-1s 23.27 0.6874 0.2807 0.2185 113.31 5.0679 71.28 0.6346 0.7235
+RealSR Ours-1s 24.81 0.7172 0.2743 0.2104 114.45 5.1298 71.19 0.6347 0.7160
 StableSR-200s 23.27 0.5722 0.3111 0.2046 24.95 4.7737 65.78 0.6164 0.6753
@@
 AddSR-4s 22.16 0.5280 0.4053 0.2360 35.41 5.2584 70.99 0.6596 0.7593
-DIV2K-Val Ours-1s 22.12 0.5595 0.2704 0.1893 23.55 4.3730 71.38 0.6136 0.7273
+DIV2K-Val Ours-1s 23.02 0.5808 0.2673 0.1821 29.16 4.3244 71.69 0.6192 0.7416
 Qualitative Comparisons.
@@
 Our approach has garnered a 69.2% user satisfaction rating.
-Table 4: Ablation study of Target Score Matching loss and Distribution-Aware Sampling Module.
-Datasets Method NIQE $\downarrow$ MUSIQ $\uparrow$ MANIQA $\uparrow$ CLIPIQA $\uparrow$
-DRealSR w/o TSM 5.6268 63.90 0.5749 0.6958
-w/o DASM 6.0368 63.56 0.5812 0.7123
-Full 5.5044 66.66 0.5850 0.7345
-RealSR w/o TSM 5.2728 68.56 0.6338 0.6987
-w/o DASM 5.5638 69.05 0.6273 0.7031
-Full 5.0679 71.28 0.6346 0.7235
 4.4 Ablation Study
-To validate the effectiveness of TSM loss and DASM, we conducted ablation studies by removing them separately in our experiments.
-We selected no-reference metrics for comparison, as these are critical for assessing image quality.
+Effectiveness of TSM and DASM.
+To validate the effectiveness of TSM loss and DASM, we conduct ablation studies by removing them separately in our experiments.
+We select LPIPS, DISTS, MUSIQ, MANIQA and CLIPIQA for comparison, as these metrics are critical for image quality assessment.
+Additionally, we select FID as the metric to evaluate distribution similarity.
 The results are presented in Tab.
 4.
-We draw the following conclusions: (1) The absence of TSM loss and DASM affects the no-reference metrics NIQE, MUSIQ, MANIQA, and CLIPIQA.
-(2) The lack of TSM leads to a significant decrease in MUSIQ and CLIPIQA metrics, possibly due to the unreliable directions in VSD causing unrealistic generations.
-(3) The absence of DASM results in a decline in NIQE, MUSIQ and CLIPIQA, possibly due to the suboptimal optimization of details.
+We draw the following conclusions: (1) The absence of TSM loss and DASM negatively impacts performance across both reference-based (LPIPS, DISTS) and non-reference-based metrics (MUSIQ, MANIQA and CLIPIQA).
+The FID metric is also adversely affected, suggesting a decline in distribution fidelity.
+(2) The lack of TSM leads to a significant decrease in LPIPS, DISTS, MUSIQ, and CLIPIQA metrics, possibly due to the unreliable directions in VSD causing unrealistic generations.
+(3) The absence of DASM results in a decline in FID, MUSIQ and CLIPIQA, possibly due to the suboptimal optimization of details.
+Table 4: Ablation study of Target Score Matching loss and Distribution-Aware Sampling Module.
+Datasets Method LPIPS $\downarrow$ DISTS $\downarrow$ FID $\downarrow$ MUSIQ $\uparrow$ MANIQA $\uparrow$ CLIPIQA $\uparrow$
+DRealSR w/o TSM 0.3176 0.2327 139.67 63.90 0.5749 0.6958
+w/o DASM 0.3097 0.2311 143.97 63.56 0.5812 0.7123
+Full 0.2967 0.2136 134.98 66.62 0.5874 0.7344
+RealSR w/o TSM 0.2934 0.2397 117.15 68.56 0.6338 0.6987
+w/o DASM 0.2873 0.2273 123.51 69.05 0.6273 0.7031
+Full 0.2743 0.2104 114.45 71.19 0.6347 0.7160
+Base model for fairer comparison.
+To validate the effectiveness of our method across different versions of SD models, we conduct additional experiments, as shown in Tab.
+5, including SD2-base and SD2.1-base models.
+The performance is evaluated on the DRealSR test dataset [52].
+It demonstrates superior performance compared to other one-step SR methods, including OSEDiff [53] and AddSR [55].
+Specifically, our SD2-base version model outperforms the single-step AddSR across all perceptual reference metrics and no-reference metrics, particularly excelling in NIQE[64], MUSIQ[22], and CLIPIQA[44], significantly surpassing the performance of one step AddSR.
+Our SD2.1-base model demonstrates comparable performance to OSEDiff and surpasses it across various metrics, with notable improvements in NIQE and CLIPIQA.
+Table 5: Same base model for fairer comparison to validate the effectiveness of TSD-SR.
+[…]
+AddSR SD2-base 0.3196 0.2242 6.9321 60.85 0.6188
+Ours SD2-base 0.3040 0.2234 6.2202 65.14 0.6935
+OSEDiff SD2.1-base 0.2968 0.2162 6.4471 64.69 0.6958
+Ours SD2.1-base 0.2943 0.2115 5.7934 65.41 0.7109
+Parameters $N$ and $s$ in DASM.
+We provide performance comparisons for different combinations of $N$ and $s$ in Tab.
+6.
+The performance is evaluated on the DRealSR test dataset.
+$N$ is set to 4, and $s$ is set to 50 in our setting (bold in the table).
+Larger or smaller $N$ will degrade performance, possibly because it is related to regularization strength.
+We prefer to use a smaller $N$ because DASM is computationally time-consuming.
+Therefore, after carefully balancing training duration and model performance, we select $N$=4 as the final value.
+Small $s$ will have similar performance, but the image quality will be compromised when setting large $s$.
+Our experimental results indicate that selecting $s$ within the range of 25-75 may yield better performance.
+Table 6: Ablation studies for hyperparameter $N$ and $s$.
+[…]
+2 50 0.3104 0.2327 137.64 64.49 0.5717 0.7118
+4 50 0.2967 0.2136 134.98 66.62 0.5874 0.7344
+8 50 0.3421 0.2633 151.64 65.73 0.5875 0.7227
+4 25 0.3063 0.2201 130.74 66.19 0.5828 0.7269
+4 100 0.3176 0.2230 135.54 65.23 0.5782 0.7024
 5 Conclusion and Limitation
@@
 In the future, we plan to employ pruning or quantization methods to compress the model parameters, striving for a lightweight and efficient Real-ISR model.
+6 Acknowledgment
+This work was supported by National Natural Science Foundation of China (62293554, U2336212), Zhejiang Provincial Natural Science Foundation of China (LD24F020007), National Key R&D Program of China (SQ2023AAA01005), Ningbo Innovation “Yongjiang 2035” Key Research and Development Programme (2024Z292), and Young Elite Scientists Sponsorship Program by CAST (2023QNRC001).
 Supplementary Material
+In this Supplementary Material, we provide additional details, including the comparison with GAN-based methods in Appendix A, more visual comparisons in Appendix B, comparisons of full-reference metrics and human preference in Appendix C, theory of Target Score Matching in Appendix D and algorithm in Appendix E.
+We conduct these additional comparisons and analyses to validate the effectiveness of TSD-SR.
 Appendix A Comparison with GAN-based Methods
 We compared our method with GAN-based approaches in Tab.
-5.
+7.
 While the GAN methods show advantages in full-reference metrics such as PSNR and SSIM, our model outperforms GAN-based methods on all no-reference metrics.
-Some research has found the limitations of full-reference methods in the field of image super-resolution [60, 54].
-Full-reference metrics often assign low scores to visually appealing results because these results produce more details, some of which may not align with the ground truth.
+Some researchers have found the limitations of PSNR and SSIM in the field of image super-resolution [61, 55].
+The effectiveness of PSNR and SSIM in assessing image fidelity in complex degradation scenarios remains debatable, as pixel-level misalignment often arises when restoring severely degraded images.
 However, no-reference metrics assess image quality based on the individual image, without the need to forcibly align with the ground truth.
 Therefore, in more complex and realistic degradation scenarios, no-reference metrics may be more suitable for evaluating the results of image super-resolution.
-In Appendix C, we discuss the comparison between full-reference metrics and human preferences, and in the Fig.
+In Appendix C, we further provide the visual comparison between full-reference metrics and human preferences, and in Fig.
 9, we present a visual comparison with GAN-based methods.
@@
 FeMASR 26.87 0.7569 0.3156 0.2238 157.72 5.9067 53.70 0.4413 0.5633
-DRealSR Ours 26.01 0.7174 0.3092 0.2216 130.83 5.5044 66.66 0.5850 0.7345
+DRealSR Ours 27.77 0.7559 0.2967 0.2136 134.98 5.9131 66.62 0.5874 0.7344
 BSRGAN 26.38 0.7651 0.2656 0.2123 141.24 5.6431 63.28 0.5425 0.5114
@@
 FeMASR 25.06 0.7356 0.2936 0.2285 141.01 5.7696 59.05 0.4872 0.5405
-RealSR Ours 23.27 0.6874 0.2807 0.2185 113.31 5.0679 71.28 0.6346 0.7235
+RealSR Ours 24.81 0.7172 0.2743 0.2104 114.45 5.1298 71.19 0.6347 0.7160
 BSRGAN 24.58 0.6269 0.3502 0.2280 49.55 4.7501 61.68 0.4979 0.5386
@@
 FeMASR 22.45 0.5858 0.3370 0.2205 41.97 4.8679 57.94 0.4787 0.5769
-DIV2K-Val Ours 22.12 0.5595 0.2704 0.1893 23.55 4.3730 71.38 0.6136 0.7273
+DIV2K-Val Ours 23.02 0.5808 0.2673 0.1821 29.16 4.3244 71.69 0.6192 0.7416
 Figure 9: Qualitative comparisons between TSD-SR and GAN-based Real-ISR methods. Please zoom in for a better view.
@@
 In Figs.
-10, 11 and 12, we provide more visual caparisons with other diffusion-based methods.
-Numerous examples demonstrate the robust restoration capabilities of TSD-SR and the high quality of the restored images.
+10, 11 and 12, we provide more visual comparisons with other diffusion-based methods.
+Numerous examples demonstrate the robust restoration capabilities of TSD-SR and the high quality of restored images.
 Figure 10: Qualitative comparisons between TSD-SR and different diffusion-based methods.
@@
 Appendix C Comparisons of Full-reference Metrics and Human Preference
-We present additional comparative experiments to demonstrate the limitations of full-reference metrics in Figure 13.
-It can be observed that GAN-based methods with advantages in PSNR and SSIM often result in blurred details and textures.
-In contrast, our results offer better visual effects, yet they do not hold an advantage in terms of PSNR and SSIM metrics.
-This phenomenon has also been discussed in other research works [3, 67, 32], with the majority of studies suggesting that for models with strong generation and restoration capabilities, these full-reference metrics do not adequately reflect image quality.
-We anticipate the development of better full-reference metrics in the future to assess advanced Real-ISR methods.
+We present additional comparative experiments in Figure 13 to demonstrate that PSNR and SSIM may have limitations in assessing image fidelity under complex degradation scenarios.
+It can be observed that GAN-based methods with higher PSNR and SSIM produce over-smooth or broken textures, raising concerns about their realism and fidelity.
+While our approach trades off PSNR and SSIM for natural detail restoration, it achieves enhanced realism and broader perceptual acceptance (Our additional user study reveals that 90.28% of participants prefer ours instead of high PSNR and SSIM methods.).
 Figure 13: Comparisons between full-reference metric assessments and human visual preference. Despite scoring lower on full-reference metrics, TSD-SR generates images that align with human preference.
@@
 The effectiveness of the method can be proven by starting from the KL divergence.
-We can use a Sticking-the-Landing style gradient by thinking of ${\epsilon}_{\psi}(\mathbf{z}_{t};c_{y},t)$ as a control variate for $\hat{\epsilon}$ For detailed proof, refer to Appendix 4 of DreamFusion [36].
+We can use a Sticking-the-Landing [39] style gradient by thinking of ${\epsilon}_{\psi}(\mathbf{z}_{t};c_{y},t)$ as a control variate for $\hat{\epsilon}$.
+For detailed proof, refer to Appendix 4 of DreamFusion [36].
 It demonstrates that the gradient of this loss yields the same updates as optimizing the training loss $\mathcal{L}_{\mathrm{MSE}}$ Eq.
@@
 Appendix E Algorithm
-Algorithm 1 details our TSD-SR training procedure. We use classifier-free guidance (cfg) for the Teacher Model and the LoRA Model. The cfg value is set to 7.5.
+Algorithm 1 details our TSD-SR training procedure. We use classifier-free guidance (cfg) for the Teacher Model and the LoRA Model. The cfg weight is set to 7.5.
 Input: $\mathcal{D}$ = $\{x_{L},x_{H},c_{y}\}$ , pre-trained Teacher Diffusion Model including VAE encoder $E_{\psi}$, denoising network $\epsilon_{\psi}$ and VAE decoder $D_{\psi}$, the number of iterations $N$ and step size $s$ of DASM.

### v4

@@
 Abstract
-Pre-trained text-to-image diffusion models are increasingly applied to real-world image super-resolution (Real-ISR) task.
+Pre-trained text-to-image diffusion models are increasingly applied to real-world image super-resolution (Real-ISR) tasks.
 Given the iterative refinement nature of diffusion models, most existing approaches are computationally expensive.
-While methods such as SinSR and OSEDiff have emerged to condense inference steps via distillation, their performance in image restoration or details recovery is not satisfied.
+While methods such as SinSR and OSEDiff have emerged to condense inference steps via distillation, their performance in image restoration or details recovery is not satisfactory.
 To address this, we propose TSD-SR, a novel distillation framework specifically designed for real-world image super-resolution, aiming to construct an efficient and effective one-step model.
@@
 40 times faster than SeeSR) compared to the past Real-ISR approaches based on pre-trained diffusion priors.
-Our code will be released at https://github.com/Microtreei/TSD-SR.
+Our code is released at https://github.com/Microtreei/TSD-SR.
 1 Introduction
-Image super-resolution (ISR) [9, 8, 23, 26] aims to transform low-quality (LQ) images, which have experienced noise or blur, into clear high-quality (HQ) images.
-Differing from traditional ISR [6, 67], which assumes a known degradation process, real-world image super-resolution (Real-ISR) [47, 63] is designed to enhance real-world images that have suffered from complex and unknown degradations, thereby offering greater practical utility.
-Generative models, particularly Generative Adversarial Networks (GANs) [11, 33, 37] and Diffusion Models (DMs) [42, 17, 40], have shown remarkable power in handling Real-ISR task.
-GAN-based methods leverage adversarial training, toggling between the generator and discriminator to produce realistic images.
-While GANs are capable of one-step inference, they are commonly hampered by issues like mode collapse and training instability [2].
-Recently, Diffusion Models (DMs) have shown impressive performance in the realm of image generation [48, 21].
-Their robust priors empower them to produce more realistic images with richer details than GAN-based methods [40, 42].
-Some researchers [29, 61, 54, 57] have successfully leveraged pre-trained DMs for Real-ISR task.
+Image super-resolution (ISR) [9, 8, 23, 26] aims to transform low-quality (LQ) images, which have been degraded by noise or blur, into clear high-quality (HQ) images.
+Unlike traditional ISR methods [6, 67], which assume a known degradation process, real-world image super-resolution (Real-ISR) [47, 63] focuses on enhancing images affected by complex and unknown degradations, thereby offering greater practical utility.
+Generative models, particularly Generative Adversarial Networks (GANs) [11, 33, 37] and Diffusion Models (DMs) [42, 17, 40], have demonstrated remarkable capabilities in tackling Real-ISR tasks.
+GAN-based methods utilize adversarial training by alternately optimizing a generator and a discriminator to produce realistic images.
+While GANs support one-step inference, they are often hindered by challenges such as mode collapse and training instability [2].
+Recently, Diffusion Models (DMs) have demonstrated impressive performance in image generation [48, 21].
+Their strong priors enable them to produce more realistic images with richer details compared to GAN-based methods [40, 42].
+Some researchers [29, 61, 54, 57] have successfully leveraged pre-trained DMs for Real-ISR tasks.
 However, due to the iterative denoising nature of diffusion models [17], the Real-ISR process is computationally expensive.
@@
 To achieve an efficient and one-step network akin to GANs, several pioneering methods that condense the iterations of diffusion models through distillation [15, 58, 12, 18] have been proposed [53, 55, 49].
-Among these works, OSEDiff [53] introduced the Variational Score Distillation (VSD) loss [51] to Real-ISR task, achieving state-of-the-art (SOTA) one-step performance by leveraging prior knowledge from pre-trained models.
+Among these works, OSEDiff [53] introduced the Variational Score Distillation (VSD) loss [51] to Real-ISR tasks, achieving state-of-the-art (SOTA) one-step performance by leveraging prior knowledge from pre-trained models.
 Despite these advancements, our investigation has revealed two critical limitations associated with VSD in Real-ISR applications.
 (1) Unreliable gradient direction.
-VSD relies on a Teacher Model to provide a “true gradient direction.” However, this guidance proves unreliable in scenarios where initial ISR outputs are suboptimal.
+VSD relies on a Teacher Model to provide a “true gradient direction.” However, this guidance is proven unreliable in scenarios where initial ISR outputs are suboptimal.
 (2) Insufficient detail recovery.
 The VSD loss exhibits notable variation across different timesteps, and the uniform sampling strategy for $t$ poses challenges in aligning the score function with detailed texture recovery requirements.
-These findings highlight the need for enhanced approaches to address these issues effectively.
+These findings underscore the need for more effective approaches to address these issues.
 In this paper, we propose a novel method called TSD-SR to distill a multi-step Text-to-Image (T2I) DMs [38, 40, 10] into an effective one-step diffusion model tailored for the Real-ISR task.
 Specifically, TSD-SR consists of two components: Target Score Distillation (TSD) and Distribution-Aware Sampling Module (DASM).
-TSD incorporates our new proposed Target Score Matching (TSM) loss to compensate for the limitation of VSD loss.
-This significant score loss leverages HQ data to provide a reliable optimization trajectory for the distillation process.
-It effectively reduces the visual artifacts caused by deviant predictions from the Teacher Model.
-DASM is crafted to bolster detail recovery by strategically sampling low-noise samples that are distribution-based during training.
-This approach effectively allocates more optimization to early timesteps in a single iteration, enhancing the recovery of details.
-Experiments on popular benchmarks demonstrate that TSD-SR achieves superior restoration performance (most of the metrics perform the best) and high efficiency (the fastest inference speed, 40 times faster than SeeSR) compared to the state-of-the-art Real-ISR methods based on pre-trained DMs, while requiring only a single inference step.
+TSD incorporates our newly proposed Target Score Matching (TSM) loss to compensate for the limitations of the VSD loss.
+This significant score loss leverages HQ data to provide a reliable optimization trajectory during distillation, effectively reducing visual artifacts caused by deviant predictions from the Teacher Model.
+DASM is designed to enhance detail recovery by strategically sampling low-noise samples that are distribution-based during training.
+This approach effectively allocates more optimization to early timesteps within a single iteration, thereby improving the recovery of fine details.
+Experiments on popular benchmarks demonstrate that TSD-SR achieves superior restoration performance (most of the metrics perform the best) and high efficiency (the fastest inference speed, 40 times faster than SeeSR) compared to the state-of-the-art Real-ISR methods based on pre-trained DMs, while requiring only a single inference step, as shown in (Fig.
+1).
 Our main contribution can be summarized as threefold:
   • We propose a novel method called TSD-SR to achieve one-step DMs distillation for the Real-ISR task.
-  • We introduce Target Score Distillation (TSD) to provide a reliable gradient to enhance the realism of the Real-ISR model’s outputs.
+  • We introduce Target Score Distillation (TSD) to provide reliable gradients that enhance the realism of outputs from Real-ISR methods.
   • We design a Distribution-Aware Sampling Module (DASM) specifically tailored to enhance the capability of detail restoration.
@@
 OSEDiff [53] directly uses LQ images as the beginning of the diffusion process, and employs VSD loss [51] as a regularization technique to condense a multi-step pre-trained T2I model into a one-step Real-ISR model.
-However, due to the incorporation of alternating training strategies, OSEDiff may initially tend towards unreliable optimization directions, which may lead to the visual artifacts.
+However, due to the incorporation of alternating training strategies, OSEDiff may initially tend towards unreliable optimization directions, which may lead to visual artifacts.
 3 Methodology
 Figure 2: Pipeline overview.
-We train one-step Student Model $G_{\theta}$ to map the low-quality image $x_{L}$ into a more realistic one.
-The noisy latent $\boldsymbol{\hat{z}_{t}}$ and $\boldsymbol{z_{t}}$ sampled by DASM (Details can be found in Fig.
-6.) will be fed into both the pre-trained Teacher and LoRA Model to produce the Variational Score loss.
-Subsequently, two Models’ predictions on $\boldsymbol{z_{t}}$ yield Target Score loss.
-Their weighted forms, namely TSD ( red flow), in conjunction with pixel-space reconstruction loss ( green flow), are leveraged to update the Student Model.
-After updating the Student Model, we employ the diffusion loss ( blue flow) to alternately update the LoRA Model.
+We train a one-step Student Model $G_{\theta}$ to transform the low-quality image $x_{L}$ into a more realistic one.
+The noisy latent $\boldsymbol{\hat{z}_{t}}$ sampled by DASM (Details can be found in Fig.
+6.) will be fed into both the pre-trained Teacher and the LoRA Model to produce the Variational Score Loss.
+Subsequently, the Teacher’s predictions on $\boldsymbol{\hat{z}_{t}}$ and $\boldsymbol{z_{t}}$ yield the Target Score Loss.
+Their weighted forms, namely TSD ( red flow), along with the pixel-space reconstruction loss ( green flow), are leveraged to update the Student Model $G_{\theta}$ After updating the Student Model, we employ the diffusion loss ( blue flow) to update the LoRA Model.
 3.1 Preliminaries
 Problem Formulation.
-ISR problem aims to reconstruct a HQ image ${x}_{H}$ from a LQ input $x_{L}$ by training a parameterized ISR model $G_{\theta}$ on a dataset $\mathcal{D}=\{(x_{L},x_{H})_{i=1}^{N}\}$, where $N$ represents the number of image pairs.
-It is mathematically formulated as minimizing the following objective:
+The ISR problem aims to reconstruct a HQ image ${x}_{H}$ from an LQ input $x_{L}$ by training a parameterized ISR model $G_{\theta}$ on a dataset $\mathcal{D}=\{(x_{L},x_{H})_{i=1}^{N}\}$, where $N$ represents the number of image pairs.
+Formally, this problem can be formulated as minimizing the following objective:
 $\theta^{*}=\arg\min_{\theta}\mathbb{E}_{(x_{L},x_{H})\sim\mathcal{D}}[\mathcal{L}_{Rec}(G_{\theta}(x_{L}),x_{H})$ (1)
 $+\lambda\mathcal{L}_{Reg}(q_{\theta}(\hat{x}_{H}),p(x_{H}))]$
-Here, $\mathcal{L}_{Rec}$ denotes the reconstruction loss, commonly measured using distance metrics like $L_{2}$ or $LPIPS$ [66].
-The regularization term $\mathcal{L}_{Reg}$ enhances the realism and generalization of the ISR model’s outputs.
-This objective can be understood as aligning the ISR output $\hat{x}_{H}$’s distribution, $q_{\theta}(\hat{x}_{H})$, with the true distribution $p(x_{H})$ by minimizing the KL-divergence [25]:
+Here, $\mathcal{L}_{Rec}$ denotes the reconstruction loss, commonly measured by distance metrics such as $L_{2}$ or $LPIPS$ [66].
+The regularization term $\mathcal{L}_{Reg}$ improves the realism and generalization of the output of the ISR model.
+This objective can be understood as aligning the ISR output $\hat{x}_{H}$’s distribution, $q_{\theta}(\hat{x}_{H})$, with the high-quality data $x_{H}$’s distribution $p(x_{H})$ by minimizing the KL-divergence [25]:
 $\min_{\theta}\mathcal{D}_{\mathrm{KL}}\left(q_{\theta}(\hat{x}_{H})\|p(x_{H})\right)$ (2)
 While several studies [47, 63, 46] have employed adversarial loss to optimize this objective, they often encounter issues like mode collapse and training instability.
-Recent work [53] achieved the state-of-the-art results using Variational Score Distillation (VSD) as the regularization loss to minimizing this objective, which inspires our research.
+Recent work [53] achieved state-of-the-art results using Variational Score Distillation (VSD) as the regularization loss to minimize this objective, which inspires our research.
 Variational Score Distillation.
-Variational Score Distillation (VSD) [51] was initially introduced for text-to-3D generation by distilling a pre-trained text-to-image diffusion model to optimize a single 3D representation [36].
+Variational Score Distillation (VSD) [51] was initially introduced for text-to-3D generation, by distilling a pre-trained text-to-image diffusion model to optimize a single 3D representation [36].
 In the VSD framework, a pre-trained diffusion model, represented as $\boldsymbol{\epsilon}_{\psi}$, and its trainable (LoRA [19]) replica $\boldsymbol{\epsilon}_{\phi}$, are used to regularize the generator network $G_{\theta}$.
@@
 As depicted in Fig.
-2, our goal is to distill a given pre-trained T2I DMs into a fast one-step Student Model $G_{\theta}$ with Teacher Model $\boldsymbol{\epsilon}_{\psi}$ and trainable LoRA Model $\boldsymbol{\epsilon}_{\phi}$.
-We denote the latent outputs of the distilled model as $\boldsymbol{\hat{z}_{0}}$, and HQ latent representations as $\boldsymbol{z_{0}}$.
-Both $\boldsymbol{\hat{z}_{0}}$ and $\boldsymbol{z_{0}}$ will pass through our Distribution-Aware Sampling Module (DASM) to obtain distribution-based samples $\boldsymbol{\hat{z}_{t}}$ and $\boldsymbol{z_{t}}$ (Sec.
+2, our goal is to distill a given pre-trained T2I DM into a fast one-step Student Model $G_{\theta}$, using the Teacher Model $\boldsymbol{\epsilon}_{\psi}$ and the trainable LoRA Model $\boldsymbol{\epsilon}_{\phi}$.
+We denote the latent output of the distilled model as $\boldsymbol{\hat{z}_{0}}$, and the HQ latent representation as $\boldsymbol{z_{0}}$.
+Both $\boldsymbol{\hat{z}_{0}}$ and $\boldsymbol{z_{0}}$ are passed through our Distribution-Aware Sampling Module (DASM) to obtain distribution-based samples $\boldsymbol{\hat{z}_{t}}$ and $\boldsymbol{z_{t}}$ (Sec.
 3.4).
-We train $G_{\theta}$ by minimizing the two losses: a reconstruction loss in pixel space to compare the model outputs against the ground truth, and a regularization loss (from Target Score Distillation) to enhances the realism (Sec.
+We train $G_{\theta}$ by minimizing the two losses: a reconstruction loss in pixel space to compare the model outputs against the ground truth, and a regularization loss (from Target Score Distillation) to enhance the realism (Sec.
 3.3).
-After updating the Student Model, we update the LoRA Model with diffusion loss.
+After updating the Student Model, we update the LoRA Model with the diffusion loss.
 Finally, in Sec.
@@
 3.3 Target Score Distillation
-Similar to [53], we introduce VSD loss into our work as a regularization loss to enhance the realism and generalization of the $G_{\theta}$ outputs.
+Similar to [53], we introduce VSD loss into our work as a regularization term to enhance the realism and generalization of the $G_{\theta}$’s outputs.
 Upon reviewing VSD Eq.
-3, $\boldsymbol{\epsilon}_{\phi}(\boldsymbol{\hat{z}_{t}};t,c_{y})$ represents the gradient direction of the optimization on $G_{\theta}$ noisy outputs $\boldsymbol{\hat{z}_{t}}$, whereas $\boldsymbol{\epsilon}_{\psi}(\boldsymbol{\hat{z}_{t}};t,c_{y})$ corresponds to the gradient direction on noisy HQ latent representations $\boldsymbol{z_{t}}$.
-The overarching goal of model optimization is to harmonize the predictions from $\boldsymbol{\hat{z}_{t}}$ with those from $\boldsymbol{z_{t}}$.
-Nonetheless, this strategy encounters hurdles, especially in the early training phase: the quality of synthetic latents $\boldsymbol{\hat{z}_{t}}$ is not high enough for Teacher Model to give a precise prediction.
+3, $\boldsymbol{\epsilon}_{\phi}(\boldsymbol{\hat{z}_{t}};t,c_{y})$ represents the current estimated gradient direction for $G_{\theta}$’s noisy outputs $\boldsymbol{\hat{z}_{t}}$, whereas $\boldsymbol{\epsilon}_{\psi}(\boldsymbol{\hat{z}_{t}};t,c_{y})$ corresponds to the ideal gradient direction guiding towards more realistic outputs.
+The overarching goal of model optimization is to align the suboptimal gradient direction with the superior direction based on pre-trained priors, thus facilitating the optimization of the Student distribution toward that of the Teacher.
+However, this strategy encounters hurdles, especially in the early training phase: the quality of synthetic latent $\boldsymbol{\hat{z}_{t}}$ is not high enough for the Teacher Model to provide a precise prediction.
 As illustrated in Fig.
-3, the Teacher Model finds it challenging to accurately predict the optimization direction for low-quality synthetic latents, with the cosine similarity to the ideal optimization direction only $0.2$ ($0.88$ on HQ latents).
-This problem can lead to severe visual artifacts, as evident in Fig.
+3, the Teacher Model struggles to accurately predict the optimization direction for low-quality synthetic latent $\boldsymbol{\hat{z}_{t}}$ in the early stage, as indicated by a cosine similarity of only $0.2$ to the ideal direction, compared to $0.88$ for high-quality latent $\boldsymbol{z_{t}}$.
+This problem can lead to severe visual artifacts, as is evident in Fig.
 4(a).
 Figure 3: A visual comparison of the gradient direction.
-We set the timestep $t$ to 100 and calculated the cosine similarity between the predicted directions from the Teacher Model and the true direction (towards the HQ data).
+We set the timestep $t$ to 100 and calculated the cosine similarity between the prediction directions from the Teacher Model and the true direction (towards the HQ data).
 The prediction direction for $\boldsymbol{z_{t}}$ closely matches the true direction, but not for $\boldsymbol{\hat{z}_{t}}$, suggesting that suboptimal samples may lead to directional deviations.
@@
 (a) The naive method introduces fake textures and fails to recover fine details.
-(b) MSE leads to over-smoothed generation result, lacking high-frequency information.
+(b) MSE leads to over-smoothed generation results, lacking high-frequency information.
 (c) Our method offers the superior visual effects and fine textures.
-A straightforward remedial measure is to employ a mean squared error (MSE) loss to align the synthetic latents with the ideal inputs of the Teacher Model, which are derived from HQ latents.
+A straightforward remedial measure is to employ a mean squared error (MSE) loss to align the synthetic latent with the ideal inputs of the Teacher Model, which are derived from the HQ latent.
 However, as shown in Fig.
 4(b), this approach has been observed to lead to over-smoothed results [13].
-Our strategy, instead, is to align the predictions made by the Teacher Model on both synthetic and HQ latents, thereby encouraging greater consistency between them.
+Our strategy, instead, is to align the predictions made by the Teacher Model on both synthetic and HQ latent, thereby encouraging greater consistency between them.
 The core idea is that for samples drawn from the same distribution, the real scores predicted by the Teacher Model should be close to each other.
 We refer to this approach as Target Score Matching (TSM):
-$\begin{split}&\nabla_{\theta}\mathcal{L}_{\mathrm{TSM}}(\mathbf{\hat{z}},\mathbf{z},c_{y})\\ &=\mathbb{E}_{t,\epsilon}\left[w(t)({\epsilon}_{\psi}(\mathbf{\hat{z}}_{t};t,c_{y})-{\epsilon}_{\psi}(\mathbf{z}_{t};t,c_{y}))\frac{\partial\mathbf{\hat{z}}}{\partial\theta}\right]\end{split}$ (4)
+$\begin{split}&\nabla_{\theta}\mathcal{L}_{\mathrm{TSM}}(\boldsymbol{\hat{z}},\boldsymbol{z},c_{y})\\ &=\mathbb{E}_{t,\epsilon}\left[w(t)({\epsilon}_{\psi}(\boldsymbol{\hat{z}}_{t};t,c_{y})-{\epsilon}_{\psi}(\boldsymbol{z}_{t};t,c_{y}))\frac{\partial\boldsymbol{\hat{z}}}{\partial\theta}\right]\end{split}$ (4)
 where the expectation of the gradient is computed across all diffusion timesteps $t\in\{1,\cdots,T\}$ and $\epsilon\sim\mathcal{N}(0,I)$.
 Equation 4 encapsulates the optimization loss for our Target Score Matching.
-Upon examining it in conjunction with the Eq.
-3, we notice that VSD utilizes the prediction residual between a Teacher and a LoRA Model to drive gradient backpropagation.
-Conversely, our TSM employs the synthetic and the HQ data to produce the gradients.
-By blending these two strategies with hyperparameter weighting $\lambda$ and $1-\lambda$, we can construct our ultimate optimization loss, effectively unifying the strengths of both approaches to optimize the training process Eq.
-5.
-$\begin{split}\nabla_{\theta}\mathcal{L}_{\mathrm{TSD}}(\mathbf{\hat{z}},\mathbf{z},c_{y})=\mathbb{E}_{t,\epsilon}\Big{[}w(t)[{\epsilon}_{\psi}(\mathbf{\hat{z}}_{t};t,c_{y})-\\ {\epsilon}_{\psi}(\mathbf{z}_{t};t,c_{y})+\lambda({\epsilon}_{\psi}(\mathbf{z}_{t};t,c_{y})-{\epsilon}_{\phi}(\mathbf{\hat{z}}_{t};t,c_{y}))]\frac{\partial\mathbf{\hat{z}}}{\partial\theta}\Big{]}\end{split}$ (5)
-where $w(t)$ is time-aware weighting function tailored for Real-ISR.
-Other symbols are in accordance with those previously mentioned.
-By introducing the prediction of pre-trained diffusion model on HQ latents, we have circumvented the issue of the model falling into the visual artifacts or over-smoothed problem, as illustrated in Fig.
+Upon examining it in conjunction with Eq.
+3, we notice that VSD utilizes the prediction residual between the Teacher and the LoRA Model to drive gradient backpropagation.
+Similarly, our TSM employs the synthetic and the HQ data to produce the gradients.
+By blending these two strategies with hyperparameter weights $\lambda$ and $1-\lambda$, we construct a combined optimization loss that effectively unifies the strengths of both approaches, as formulated in Eq.
+5, to guide the training process.
+$\begin{split}\nabla_{\theta}\mathcal{L}_{\mathrm{TSD}}(\boldsymbol{\hat{z}},\boldsymbol{z},c_{y})=\mathbb{E}_{t,\epsilon}\Big{[}w(t)[{\epsilon}_{\psi}(\boldsymbol{\hat{z}}_{t};t,c_{y})-\\ {\epsilon}_{\psi}(\boldsymbol{z}_{t};t,c_{y})+\lambda({\epsilon}_{\psi}(\boldsymbol{z}_{t};t,c_{y})-{\epsilon}_{\phi}(\boldsymbol{\hat{z}}_{t};t,c_{y}))]\frac{\partial\boldsymbol{\hat{z}}}{\partial\theta}\Big{]}\end{split}$ (5)
+where $w(t)$ is a time-aware weighting function tailored for Real-ISR.
+Other symbols are consistent with those previously defined.
+By introducing the prediction of the pre-trained diffusion model on HQ latent, we have circumvented the issue of the model falling into visual artifacts or producing over-smoothed results, as illustrated in Fig.
 4(c).
 3.4 Distribution-Aware Sampling Module
-In the VSD-based framework, we need to match the score functions predicted by the Teacher Model with those by the LoRA Model at different timesteps $t\sim\{0,1,\cdots,T\}$.
-However, for the Real-ISR problem, this matching ability at different $t$ is unbalanced, as illustrated in Fig.
+In the VSD-based framework, it is necessary to match the score functions predicted by the Teacher Model and the LoRA Model across timesteps $t\in{0,1,\dots,T}$.
+However, for the Real-ISR problem, this matching performance is inconsistent across timesteps, as illustrated in Fig.
 5(a).
-This phenomenon could be attributed to our leveraging of the low-frequency (LF) priors from LQ data ,but losing guidance from high-frequency (HF) details.
-The noisy samples $\hat{z}_{t}$ possesses low-frequency priors (from LQ data).
-These priors are easily learned by the LoRA Model, leading to more similar predictions (Fig.
-5(b)) during LF restoration (Stage 1).
-However, In the Stage 2, the lack of HF details in $\hat{z}_{t}$ causes the LoRA Model to poorly perceive details, resulting in divergent predictions (Fig.
-5(c)).
-Our goal is to reduce this divergence.
+This phenomenon may be attributed to the reliance on low-frequency (LF) priors in the LQ data, while lacking guidance from high-frequency (HF) details.
+The output sample $\boldsymbol{\hat{z}_{0}}$, derived from LQ data, contains low-frequency (LF) priors that are easily captured by the LoRA Model, resulting in similar predictions during LF restoration (Stage 1), as shown in Fig.
+5(b).
+However, in Stage 2, due to the absence of high-frequency (HF) details in $\boldsymbol{\hat{z}_{0}}$, the LoRA Model struggles to reconstruct fine-grained features, leading to divergent predictions, as illustrated in Fig.
+5(c).
+To address this issue, we aim to reduce such divergence.
 Figure 5: (a) The prediction errors of the VSD loss at different timesteps.
@@
 (c) The visualization of Stage 2 prediction error.
-Existing method, for each iteration, only matches the score function at a single timestep sample $\hat{z}_{t}$, with uniform sampling strategy of time $t$.
-This leads to slow convergence or even difficulty in optimization during Stage 2, as the gradients of important timesteps are averaged out.
+Existing methods match the score function at each iteration using a single latent sample $\boldsymbol{\hat{z}_{t}}$, with the timestep $t$ drawn from a uniform distribution.
+This leads to slow convergence and even training difficulty during Stage 2, as gradients from important timesteps are diluted by uniform averaging.
 To this end, we propose our Distribution-Aware Sampling Module (DASM).
@@
 As shown in Fig.
-6, we first introduce noise into the synthetic latent representation $\boldsymbol{\hat{z}_{t}}=(1-\sigma_{t})\boldsymbol{\hat{z}_{0}}+\sigma_{t}\boldsymbol{\epsilon}$, where $\sigma$ is a weight factor and $\epsilon$ is a gaussian noise.
-Subsequently, we employ a LoRA Model to perform denoising, yielding the previous time noisy samples by Eq.
+6, we first obtain the noisy synthetic latent representation as $\boldsymbol{\hat{z}_{t}}=(1-\sigma_{t})\boldsymbol{\hat{z}_{0}}+\sigma_{t}\boldsymbol{\epsilon}$, where $\sigma_{t}$ is a weighting factor and $\epsilon$ denotes Gaussian noise.
+Subsequently, we employ a LoRA Model to perform denoising, yielding noisy samples at the previous timestep as described in Eq.
 6:
 $\boldsymbol{\hat{z}_{t-1}}=\boldsymbol{\hat{z}_{t}}+(\sigma_{t-1}-\sigma_{t})\cdot\boldsymbol{\epsilon}_{\phi}(\boldsymbol{\hat{z}_{t}};t,c_{y}),$ (6)
-where $\sigma_{t-1}$ and $\sigma_{t}$ are from the flow matching scheduler, and here LoRA Model has fit $\boldsymbol{\hat{z}_{0}}$’s distribution.
-Similarly, $\boldsymbol{z_{t-1}}$ can be obtained by denoising on the Teacher Model.
-Ultimately, in a single iteration, the gradients from noise samples along its trajectory can be accumulated to update the Student Model.
-Since the obtained samples follow the diffusion sampling trajectory and are directed towards early timestep, we effectively optimize the divergence in Stage 2.
+The parameters $\sigma_{t-1}$ and $\sigma_{t}$ are obtained from the flow matching scheduler.
+Here, the LoRA Model has learned the distribution of $\boldsymbol{\hat{z}_{0}}$.
+Similarly, $\boldsymbol{z_{t-1}}$ can be obtained by denoising using the Teacher Model.
+In a single iteration, gradients from noisy samples along the sampling trajectory can be accumulated to update the Student Model.
+Since these samples follow the diffusion sampling trajectory and are concentrated at early timesteps, this approach effectively reduces the divergence observed in Stage 2.
 Figure 6: Illustration of DASM.
-Top: The naive method that add noise to samples.
-Bottom: Our DASM leverages the priors of diffusion models to sample noise that more closely aligns with the true sampling trajectory.
-The noisy samples can all serve as inputs to the subsequent network to generate gradient backpropagation.
+Top: The naive approach that adds noise directly to the samples.
+Bottom: The proposed DASM leverages diffusion model priors to generate noisy latent that better align with the true sampling trajectory.
+These noisy samples can all serve as inputs to the downstream network, enabling effective gradient backpropagation.
 3.5 Training Objective
@@
 Student Model $G_{\theta}$. We train our Student Model with the reconstruction loss $\mathcal{L}_{Rec}$ and the regularization loss $\mathcal{L}_{Reg}$.
-For the reconstruction loss, we use $LPIPS$ loss and $MSE$ loss in the latent space:
-$\begin{split}&\mathcal{L}_{Rec}\left(G_{\theta}(\boldsymbol{x}_{L}),\boldsymbol{x}_{H}\right)\\ &=\gamma_{1}\mathcal{L}_{LPIPS}\left(G_{\theta}(\boldsymbol{x}_{L}),\boldsymbol{x}_{H}\right)+\mathcal{L}_{MSE}\left(\mathbf{z}_{t},\mathbf{\hat{z}}_{t}\right).\end{split}$ (7)
-For the regularization loss, we use our TSD loss Eq. 5. Therefore, the overall training objective for the Student Model $G_{\theta}$ is:
+For the reconstruction loss, we use the $LPIPS$ loss in the pixel space and the $MSE$ loss in the latent space:
+$\begin{split}&\mathcal{L}_{Rec}\left(G_{\theta}(\boldsymbol{x}_{L}),\boldsymbol{x}_{H}\right)\\ &=\gamma_{1}\mathcal{L}_{LPIPS}\left(G_{\theta}(\boldsymbol{x}_{L}),\boldsymbol{x}_{H}\right)+\mathcal{L}_{MSE}\left(\boldsymbol{z}_{t},\boldsymbol{\hat{z}}_{t}\right).\end{split}$ (7)
+For the regularization loss, we use our TSD loss, Eq. 5. Therefore, the overall training objective for the Student Model $G_{\theta}$ is:
 $\mathcal{L}_{Stu}=\mathcal{L}_{Rec}+\gamma_{2}\mathcal{L}_{Reg},$ (8)
@@
 As optimization progresses, we ramp up $\gamma_{1}$ from 1 to 2 while maintaining $\gamma_{2}$ at its initial value.
-LoRA Model $\boldsymbol{\epsilon}_{\phi}$. As stipulated by VSD, the fine-tuned replica $\epsilon_{\phi}$ must be trainable, with its training objective being:
+LoRA Model $\boldsymbol{\epsilon}_{\phi}$. As stipulated by VSD, the replica $\epsilon_{\phi}$ must be trainable, with its training objective being:
 $\mathcal{L}_{Diff}(\boldsymbol{{\hat{z}}},c_{y})=\mathbb{E}_{t,\boldsymbol{\epsilon}}[\left\|\boldsymbol{\epsilon}_{\phi}(\boldsymbol{\hat{z}_{t}};t,c_{y})-\boldsymbol{\epsilon}^{\prime}\right\|^{2}],$ (9)
-where $\boldsymbol{\epsilon}^{\prime}$ serves as the training target for the denoising network, representing Gaussian noise in the context of DDPM, and a gradient towards HQ for flow matching.
+where $\boldsymbol{\epsilon}^{\prime}$ serves as the training target for the denoising network, representing Gaussian noise in the context of DDPM, and a gradient towards HQ data for flow matching.
 4 Experiments
@@
 Training Datasets.
-We adopt DIV2K [1], Flickr2K [43], LSDIR [27] and first 10K face images from FFHQ [20] for training.
-We use the same degradation pipeline as Real-ESRGAN [47] to synthesize LR-HR pairs.
+For training, we utilize DIV2K [1], Flickr2K [43], LSDIR [27], and the first 10K face images from FFHQ [20].
+To synthesize LR-HR pairs, we adopt the same degradation pipeline as in Real-ESRGAN [47].
 Test Datasets.
-We evaluated our model on synthetic DIV2K-Val [1] dataset and two real-world datasets, RealSR [4] and DRealSR [52].
-The real-world datasets comprised 128x128 LQ and 512x512 HQ image pairs.
-The synthetic dataset included 3,000 image pairs, which are created by downgrading the 512x512 HQ images croped from DIV2K-Val to 128x128 LQ images using the Real-ESRGAN [47] degradation pipeline.
+We evaluate our model on the synthetic DIV2K-Val [1] dataset, as well as two real-world datasets: RealSR [4] and DRealSR [52].
+The real-world datasets consist of 128×128 low-quality (LQ) and 512×512 high-quality (HQ) image pairs.
+For the synthetic set, 3,000 pairs were generated by cropping 512×512 patches from DIV2K-Val and applying the Real-ESRGAN [47] degradation pipeline to downsample them to 128×128.
 Evaluation Metrics.
-For evaluating our method, we apply both full-reference and no-reference metrics.
-Full-reference metrics include PSNR and SSIM [50] (calculated on the Y channel in YCbCr space) for fidelity, LPIPS [66] and DISTS [7] for perceptual quality, and FID [14] for distribution comparison.
-No-reference metrics include NIQE [64], MANIQA [56], MUSIQ [22] and CLIPIQA [44].
+To evaluate our method, we employ both full-reference and no-reference metrics.
+The full-reference metrics include PSNR and SSIM [50] (computed on the Y channel of the YCbCr color space) for fidelity; LPIPS [66] and DISTS [7] for perceptual quality; and FID [14] for measuring distribution similarity.
+The no-reference metrics include NIQE [64], MANIQA [56], MUSIQ [22], and CLIPIQA [44].
 Compared Methods.
@@
 Implementation Details.
-All models mentioned are initialized from the Teacher Model (SD3 [10] in our work).
-Similar to OSEDiff [53], we only train the vae encoder and denoising network in the Student Model, frozen the vae decoder in order to preserve vae’s prior [24].
-We utilize the default prompt for the Student Model, and prompts are extracted from HQ for the Teacher and LoRA models when training.
-We use AdamW optimizer [30] with a learning rate of 5e-6 for the Student Model and 1e-6 for the LoRA Model, and set the rank of LoRA to 64 for both two models.
-During the initial training phase, we incorporate MSE loss in the latent space and exclude DASM to enhance model stability and reduce training time.
-In later stages, we remove this MSE loss as it leads to over-smoothed results, and we employ DASM to improve image restoration.
-The training process took roughly 96h, utilizing 8 NVIDIA V100 GPUs with a batch size of 16.
+All models are initialized from the Teacher Model (SD3 [10] in our work).
+Similar to OSEDiff [53], we only train the VAE encoder and the denoising network in the Student Model, freezing the VAE decoder to preserve its prior [24].
+We utilize the default prompt for the Student Model, while prompts are extracted from HQ images for the Teacher and LoRA models during training.
+We adopt the AdamW optimizer [30] with a learning rate of 5e-6 for the Student Model and 1e-6 for the LoRA Model, setting the LoRA rank to 64 for both models.
+During the initial training phase, we incorporate MSE loss in the latent space and exclude DASM to stabilize training and reduce time cost.
+In later stages, we remove the MSE loss to avoid over-smoothed results and introduce DASM to enhance restoration quality.
+The training process took approximately 96 hours, utilizing 8 NVIDIA V100 GPUs with a batch size of 16.
 Table 1: Quantitative comparison with the state-of-the-art one-step methods across both synthetic and real-world benchmarks.
@@
 Tab.
-1 show the quantitative comparison of our method with single-step diffusion models on three datasets.
-Ours achieves the best results on most evaluation metrics.
-SinSR and AddSR, as distilled versions of previous multi-step super-resolution methods, reduce inference steps but experience a corresponding decrease in performance metrics.
-OSEDiff introduces the VSD loss from 3D generation tasks into Real-ISR without fully addressing the substantial differences between the two domains.
-Therefore, the no-reference metrics for image quality are not very satisfied.
-In contrast, our proposed TSD-SR, tailored for Real-ISR, outperforms other single-step models in terms of the vast majority of key metrics.
+1 shows the quantitative comparison of our method with other single-step diffusion models on three datasets.
+Our method achieves the best results across most evaluation metrics.
+SinSR and AddSR, as distilled versions of previous multi-step super-resolution methods, significantly reduce inference steps while suffering a corresponding drop in performance.
+OSEDiff introduces the VSD loss from 3D generation tasks into the Real-ISR without fully accounting for the substantial differences between the two domains.
+As a result, its no-reference image quality metrics are not satisfactory.
+In contrast, our proposed TSD-SR, specifically designed for Real-ISR, outperforms all other single-step models in the vast majority of key metrics.
 Tab.
-2 show the quantitative comparison with multi-step models.
-We can draw the following conclusions: (1) TSD-SR demonstrates significant advantages over competing methods in terms of LPIPS, DISTS, and NIQE metrics.
-Additionally, it achieves performance that surpasses most multi-step models on FID, MUSIQ, and CLIPIQA.
-(2) DiffBIR, SeeSR, PASD, and AddSR exhibit better performance on the MANIQA metric, which may be attributed to the fact that multi-step models have more denoising iterations to produce rich details.
-(3) ResShift stands out with the highest PSNR and SSIM scores, while StableSR also shows notable performance regarding DISTS and FID metrics.
-However, both two models underperform in terms of no-reference metrics.
-Finally, we explain the lower PSNR and SSIM metrics in our experiments.
-Several works [55, 61] have found that these reconstruction metrics are not well-suited for evaluating Real-ISR tasks.
-When a model recovers better details, it leads to lower reconstruction metrics, indicating a trade-off [3, 68, 32].
-This phenomenon has also been widely discussed in other related research work [3, 68, 32, 61, 66, 45, 55].
-LPIPS [66] is proposed to overcome the limitation that PSNR and SSIM fail to align with human judgments in spatial ambiguities situation.
+2 shows the quantitative comparison with multi-step models.
+We can draw the following conclusions: (1) TSD-SR demonstrates significant advantages over competing methods in terms of LPIPS, DISTS, and NIQE.
+Additionally, it outperforms most multi-step models in FID, MUSIQ, and CLIPIQA.
+(2) DiffBIR, SeeSR, PASD, and AddSR achieve better results in terms of MANIQA, which may be attributed to the fact that multi-step models benefit from more denoising iterations to generate richer details.
+(3) ResShift stands out with the highest PSNR and SSIM scores, while StableSR also performs well in terms of DISTS and FID.
+However, both models underperform on the no-reference metrics.
+Finally, we explain the relatively lower PSNR and SSIM scores observed in our experiments.
+Several studies [55, 61] have shown that these reconstruction metrics are not well-suited for the evaluation of Real-ISR tasks.
+Models that recover more realistic or detailed textures often yield lower PSNR and SSIM scores, reflecting a fundamental trade-off between perceptual quality and pixel-wise fidelity [3, 68, 32].
+This phenomenon has also been extensively discussed in the other research work [3, 68, 32, 61, 66, 45, 55].
+LPIPS [66] is proposed to overcome the limitation that PSNR and SSIM fail to align with human judgments in spatial ambiguities situations.
 Other DMs-based SR researchers [61, 45] argue that DMs introduce superior pre-trained priors, enabling the restoration of information that traditional methods (from scratch) cannot achieve.
@@
 Fig.
-7 present visual comparisons of different Real-ISR methods.
-As shown in multi-step methods’ results, SeeSR uses degradation-aware semantic cues to leverage image generation priors, but it sometimes produces over-smooth textures.
-SUPIR demonstrates a notably robust generative capacity, however, the overproduction of extraneous details will lead to results that lack naturalism in image restoration (e.g.
-Adding excessive wrinkles at the corners of a young girl’s eyes).
-Under more realistic degradation conditions, PASD finds it challenging to restore the corresponding content, revealing a deficiency in the model’s robustness capabilities.
-Among single-step methods, SinSR introduces artifacts, possibly because its distilled pre-trained model is trained from scratch, lacking sufficient real-world priors, which results in unsatisfactory image restoration capabilities.
+7 presents visual comparisons of different Real-ISR methods.
+As shown in the results of multi-step methods, SeeSR leverages degradation-aware semantic cues to incorporate image generation priors, but it tends to produce over-smoothed textures in some cases.
+SUPIR demonstrates notably robust generative capabilities.
+However, the excessive generation of fine details can result in outputs that appear less natural (e.g., adding unnecessary wrinkles around the eyes of a young girl).
+Under more realistic degradation conditions, PASD finds it difficult to recover the appropriate content, indicating limited robustness.
+Among single-step methods, SinSR tends to produce artifacts, likely due to its base model, ResShift, being trained from scratch without adequate exposure to real-world priors, which leads to inferior image restoration quality.
 AddSR produces over-smoothed results when using its 1-step model.
-OSEDiff offers improved restoration effects compared to SinSR and AddSR, yet it may fall short in terms of authenticity and naturalness when it comes to detail recovery.
+OSEDiff demonstrates better restoration performance than SinSR and AddSR; however, it may fall short in terms of authenticity and naturalness, particularly in recovering fine details.
 In contrast, our method effectively generates rich textures and realistic details with enhanced sharpness and contrast.
@@
 Complexity Comparisons
-Table 3: Comparison of computational complexity across different DMs-based methods.
-The performance of each method is measured on an A100 GPU, using input images sized at 512 × 512 pixels.
-We have disregarded the loading time for model weights and data.
+Table 3: Comparison of computational complexity across different diffusion model-based methods.
+Performance is measured on an A100 GPU using 512×512 input images, excluding model weight and data loading time.
 StableSR DiffBIR SeeSR SUPIR PASD ResShift AddSR AddSR SinSR OSEDiff Ours
@@
 Inference Time 12.4151 7.9637 5.8167 16.8704 4.8441 0.7546 1.0199 0.5043 0.1424 0.1650 0.1362
-We assess the complexity of the state-of-the-art DM-based Real-ISR models as detailed in Tab.
-3, focusing on inference time.
-The performance of each method is benchmarked on an A100 GPU with input images sized at 512 × 512 pixels.
-We have disregarded the loading time for model weights and data, and the main computation time includes: (1) text extraction time when a text extractor is used; (2) text encoder computation time if needed; (3) VAE encoding and decoding time; (4) denoising network execution time.
-It is evident that TSD-SR has a substantial advantage in inference time compared with multi-step models.
-Specifically, TSD-SR is more than 120 times faster than SUPIR, 90 times faster than StableSR, about 50 times faster than DiffBIR, more than 40 times faster than SeeSR, more than 35 times faster than PASD, and 4 times faster than ResShift.
-When compared with existing one-step models, our model boasts the fastest inference times.
-This is because we directly denoise from LQ data and use a fixed prompt.
+We assess the computational complexity of the state-of-the-art DM-based Real-ISR methods, as detailed in Tab.
+3, with a focus on inference time.
+Each method is benchmarked on an A100 GPU using input images of size 512×512 pixels.
+We disregarded the loading time for model weights and data.
+The main computation time consists of: (1) text extraction time (if a text extractor is used); (2) text encoder computation time (if applicable); (3) VAE encoding and decoding time; and (4) denoising network execution time.
+It is evident that TSD-SR holds a substantial advantage in inference speed compared to multi-step models.
+Specifically, TSD-SR is over 120× faster than SUPIR, 90× faster than StableSR, approximately 50× faster than DiffBIR, over 40× faster than SeeSR, more than 35× faster than PASD, and 4× faster than ResShift.
+When compared with existing one-step models, our method achieves the fastest inference times.
+This advantage is attributed to directly denoising from LQ data and employing a fixed prompt.
 4.3 User Study
 We conduct a user study comparing our method with three other diffusion-based one-step super-resolution methods.
-To make the evaluation more comprehensive, we selected images from five categories—human faces, buildings, animals, vegetation and characters.
-A total of 50 participants were engaged in the voting process.
-We guided participants to evaluate the best restoration results based on the similarity to HQ image, structural similarity to the LQ image, and realism of textures and details.
-The results shown in the Fig.
-8 indicate that our method receives a 69.2% approval rate from users.
-We scored 57.60% in Animals, 70.00% in Buildings, 68.80% in Human Faces, 65.20% in Vegetation, and 84.40% in Characters, surpassing other models.
-Figure 8: The results of our user study.
-Left: Category-based user preference radar chart.
-Our model has won the most user favor across all categories.
-Right: User preference pie chart.
-Our approach has garnered a 69.2% user satisfaction rating.
+To ensure a comprehensive evaluation, we selected images from five categories—human faces, buildings, animals, vegetation, and characters.
+A total of 50 participants took part in the voting process.
+Participants were instructed to select the best restoration results based on similarity to the HQ image, structural similarity to the LQ image, and realism of textures and details.
+The results in Fig.
+8 indicate that our method received a 69.2% approval rate from users.
+Specifically, our method achieved 57.6% in Animals, 70.0% in Buildings, 68.8% in Human Faces, 65.2% in Vegetation, and 84.4% in Characters, surpassing those of other methods.
+Figure 8: Results of our user study.
+Left: Category-based user preference radar chart, showing that our model received the highest favor across all categories.
+Right: User preference pie chart, illustrating that our approach garnered a 69.2% user satisfaction rating.
 4.4 Ablation Study
 Effectiveness of TSM and DASM.
-To validate the effectiveness of TSM loss and DASM, we conduct ablation studies by removing them separately in our experiments.
-We select LPIPS, DISTS, MUSIQ, MANIQA and CLIPIQA for comparison, as these metrics are critical for image quality assessment.
-Additionally, we select FID as the metric to evaluate distribution similarity.
+To validate the effectiveness of the TSM loss and DASM, we conduct ablation studies by removing each component separately.
+We select LPIPS, DISTS, MUSIQ, MANIQA, and CLIPIQA for comparison, as these metrics are critical for image quality assessment.
+Additionally, FID is used to evaluate distribution similarity.
 The results are presented in Tab.
 4.
-We draw the following conclusions: (1) The absence of TSM loss and DASM negatively impacts performance across both reference-based (LPIPS, DISTS) and non-reference-based metrics (MUSIQ, MANIQA and CLIPIQA).
-The FID metric is also adversely affected, suggesting a decline in distribution fidelity.
-(2) The lack of TSM leads to a significant decrease in LPIPS, DISTS, MUSIQ, and CLIPIQA metrics, possibly due to the unreliable directions in VSD causing unrealistic generations.
-(3) The absence of DASM results in a decline in FID, MUSIQ and CLIPIQA, possibly due to the suboptimal optimization of details.
+From the results, we draw the following conclusions: (1) The absence of TSM loss and DASM negatively impacts performance across both reference-based metrics (LPIPS, DISTS) and no-reference metrics (MUSIQ, MANIQA, and CLIPIQA).
+The FID metric is also adversely affected, indicating a decline in distribution fidelity.
+(2) The lack of TSM leads to a significant decrease in LPIPS, DISTS, MUSIQ, and CLIPIQA, possibly due to unreliable directions in VSD leading to unrealistic generations.
+(3) The absence of DASM results in a decline in FID, MUSIQ, and CLIPIQA, possibly due to suboptimal detail optimization.
 Table 4: Ablation study of Target Score Matching loss and Distribution-Aware Sampling Module.
@@
 Base model for fairer comparison.
-To validate the effectiveness of our method across different versions of SD models, we conduct additional experiments, as shown in Tab.
-5, including SD2-base and SD2.1-base models.
+To validate the effectiveness of our method across different versions of SD models, we conduct additional experiments on SD2-base and SD2.1-base models, as shown in Tab.
+5.
 The performance is evaluated on the DRealSR test dataset [52].
-It demonstrates superior performance compared to other one-step SR methods, including OSEDiff [53] and AddSR [55].
-Specifically, our SD2-base version model outperforms the single-step AddSR across all perceptual reference metrics and no-reference metrics, particularly excelling in NIQE[64], MUSIQ[22], and CLIPIQA[44], significantly surpassing the performance of one step AddSR.
-Our SD2.1-base model demonstrates comparable performance to OSEDiff and surpasses it across various metrics, with notable improvements in NIQE and CLIPIQA.
-Table 5: Same base model for fairer comparison to validate the effectiveness of TSD-SR.
+Our method demonstrates superior performance compared to other one-step SR methods, including OSEDiff [53] and AddSR [55].
+Specifically, our SD2-base model outperforms single-step AddSR across all perceptual reference and no-reference metrics, particularly excelling in NIQE [64], MUSIQ [22], and CLIPIQA [44].
+Meanwhile, our SD2.1-base model shows comparable or better performance than OSEDiff across various metrics, with notable improvements in NIQE and CLIPIQA.
+Table 5: Fair comparison using the same base model to validate TSD-SR
 […]
@@
 Parameters $N$ and $s$ in DASM.
-We provide performance comparisons for different combinations of $N$ and $s$ in Tab.
+We compare performance under different combinations of $N$ and $s$ in Tab.
 6.
-The performance is evaluated on the DRealSR test dataset.
-$N$ is set to 4, and $s$ is set to 50 in our setting (bold in the table).
-Larger or smaller $N$ will degrade performance, possibly because it is related to regularization strength.
-We prefer to use a smaller $N$ because DASM is computationally time-consuming.
-Therefore, after carefully balancing training duration and model performance, we select $N$=4 as the final value.
-Small $s$ will have similar performance, but the image quality will be compromised when setting large $s$.
-Our experimental results indicate that selecting $s$ within the range of 25-75 may yield better performance.
+The evaluation is conducted on the DRealSR test dataset.
+In our setting, $N$ is set to 4 and $s$ to 50 (highlighted in bold in the table).
+Performance degrades when $N$ is either larger or smaller, possibly due to its effect on regularization strength.
+Since DASM is computationally expensive, we prefer a smaller $N$.
+After balancing training time and performance, we select $N=4$ as the final value.
+Smaller values of $s$ yield similar performance, while larger values degrade image quality.
+Experimental results suggest that choosing $s$ between 25 and 75 achieves better overall performance.
 Table 6: Ablation studies for hyperparameter $N$ and $s$.
@@
 5 Conclusion and Limitation
-We propose TSD-SR, an effective and one-step model based on diffusion prior for Real-ISR.
-TSD-SR utilizes the TSD to enhance the realism of images generated by the distillation model.
-And it leverages DASM to sample distribution-based samples and accumulate their gradients to enhance the recovery of details.
-Our experiments have demonstrated that TSD-SR outperforms existing one-step Real-ISR models in both performance and inference speed.
+We propose TSD-SR, an effective one-step model for Real-ISR based on diffusion priors.
+TSD-SR utilizes TSD to enhance the realism of images generated by the distillation model, and leverages DASM to sample distribution-based noisy latents and accumulate their gradients, thereby improving detail recovery.
+Our experiments demonstrate that TSD-SR outperforms existing one-step Real-ISR models in both restoration quality and inference speed.
 limitations.
-Although our model boasts excellent inference speed and restoration performance, it still has a large number of model parameters compared to the past GAN/non-diffusion approaches.
-In the future, we plan to employ pruning or quantization methods to compress the model parameters, striving for a lightweight and efficient Real-ISR model.
+Although our model achieves excellent inference speed and restoration performance, it still contains significantly more parameters compared to previous GAN- or non-diffusion-based methods.
+In future work, we plan to apply pruning or quantization techniques to compress the model, aiming to develop a lightweight and efficient Real-ISR solution.
 6 Acknowledgment
@@
 Appendix A Comparison with GAN-based Methods
-We compared our method with GAN-based approaches in Tab.
+We compare our method with GAN-based approaches in Tab.
 7.
-While the GAN methods show advantages in full-reference metrics such as PSNR and SSIM, our model outperforms GAN-based methods on all no-reference metrics.
-Some researchers have found the limitations of PSNR and SSIM in the field of image super-resolution [61, 55].
-The effectiveness of PSNR and SSIM in assessing image fidelity in complex degradation scenarios remains debatable, as pixel-level misalignment often arises when restoring severely degraded images.
-However, no-reference metrics assess image quality based on the individual image, without the need to forcibly align with the ground truth.
-Therefore, in more complex and realistic degradation scenarios, no-reference metrics may be more suitable for evaluating the results of image super-resolution.
-In Appendix C, we further provide the visual comparison between full-reference metrics and human preferences, and in Fig.
+While the GAN methods show advantages in full-reference metrics such as PSNR and SSIM, our model outperforms them across all no-reference metrics.
+Prior studies have highlighted the limitations of PSNR and SSIM for evaluating image super-resolution performance [61, 55].
+Their effectiveness in assessing image fidelity in complex degradation scenarios remains debatable, as pixel-level misalignment often arises when restoring severely degraded images.
+However, no-reference metrics evaluate image quality based solely on the individual image, without requiring alignment with the ground truth.
+Therefore, in more complex and realistic degradation scenarios, they may offer a more appropriate evaluation of super-resolution results.
+In Appendix C, we further provide a visual comparison between full-reference metrics and human preferences, and in Fig.
 9, we present a visual comparison with GAN-based methods.
-From the visualization, it can be observed that our model achieves better results in terms of texture details compared to GAN methods.
+From these visualizations, it is evident that our model produces more realistic texture details than the GAN-based approaches.
 Table 7: Quantitative comparison with GAN-based methods on both synthetic and real-world benchmarks. The best results of each metric are highlighted in red.
@@
 In Figs.
-10, 11 and 12, we provide more visual comparisons with other diffusion-based methods.
-Numerous examples demonstrate the robust restoration capabilities of TSD-SR and the high quality of restored images.
+10, 11 and 12, we provide additional visual comparisons with other diffusion-based methods.
+These examples further demonstrate the robust restoration capabilities of TSD-SR and the high quality of the restored images.
 Figure 10: Qualitative comparisons between TSD-SR and different diffusion-based methods.
@@
 Appendix C Comparisons of Full-reference Metrics and Human Preference
-We present additional comparative experiments in Figure 13 to demonstrate that PSNR and SSIM may have limitations in assessing image fidelity under complex degradation scenarios.
-It can be observed that GAN-based methods with higher PSNR and SSIM produce over-smooth or broken textures, raising concerns about their realism and fidelity.
-While our approach trades off PSNR and SSIM for natural detail restoration, it achieves enhanced realism and broader perceptual acceptance (Our additional user study reveals that 90.28% of participants prefer ours instead of high PSNR and SSIM methods.).
+We present additional comparative experiments in Figure 13 to demonstrate the limitations of PSNR and SSIM in assessing image fidelity under complex degradation scenarios.
+As observed, GAN-based methods with higher PSNR and SSIM scores tend to produce over-smoothed or fragmented textures, raising concerns about their realism and perceptual fidelity.
+In contrast, our approach sacrifices some PSNR and SSIM performance to achieve more natural detail restoration, resulting in enhanced realism and broader perceptual acceptance.
+An additional user study shows that 90.28% of participants prefer our results over those of methods with higher PSNR and SSIM scores.
 Figure 13: Comparisons between full-reference metric assessments and human visual preference. Despite scoring lower on full-reference metrics, TSD-SR generates images that align with human preference.
@@
 The core idea of Target Score Matching (TSM) is that for samples drawn from the same distribution, the real scores predicted by the Teacher Model should be close to each other.
-Thus we minimize the MSE loss between the Teacher Model’s predictions of $\mathbf{\hat{z}}_{t}$ and $\mathbf{z}_{t}$ by
-$\begin{split}&\mathcal{L}_{\mathrm{MSE}}(\mathbf{\hat{z}},\mathbf{z},c_{y})\\ &=\mathbb{E}_{t,\epsilon}\left[w(t)\|\epsilon_{\psi}(\mathbf{\hat{z}}_{t};t,c_{y})-\epsilon_{\psi}(\mathbf{z}_{t};t,c_{y})\|_{2}^{2}\right]\end{split}$ (10)
+Thus, we minimize the MSE loss between the Teacher Model’s predictions on $\boldsymbol{\hat{z}}_{t}$ and $\boldsymbol{z}_{t}$ by
+$\begin{split}&\mathcal{L}_{\mathrm{MSE}}(\boldsymbol{\hat{z}},\boldsymbol{z},c_{y})\\ &=\mathbb{E}_{t,\epsilon}\left[w(t)\|\epsilon_{\psi}(\boldsymbol{\hat{z}}_{t};t,c_{y})-\epsilon_{\psi}(\boldsymbol{z}_{t};t,c_{y})\|_{2}^{2}\right]\end{split}$ (10)
 where the expectation of the gradient is computed across all diffusion timesteps $t\in\{1,\cdots,T\}$ and $\epsilon\sim\mathcal{N}(0,I)$.
 To understand the difficulties of this approach, consider the gradient of
-$\begin{split}&\nabla_{\theta}\mathcal{L}_{\mathrm{MSE}}(\mathbf{\hat{z}},\mathbf{z},c_{y})=\mathbb{E}_{t,\epsilon}\Big{[}w(t)\cdot\underbrace{\frac{\partial{\epsilon}_{\psi}(\mathbf{\hat{z}}_{t};t,c_{y})}{\partial\mathbf{\hat{z}}_{t}}}_{\text{Diffusion Jacobian}}\\ &\underbrace{({\epsilon}_{\psi}(\mathbf{\hat{z}}_{t};t,c_{y})-{\epsilon}_{\psi}(\mathbf{z}_{t};t,c_{y}))}_{\text{Prediction Residual}}\underbrace{\frac{\partial\mathbf{\hat{z}}}{\partial\theta}}_{\text{Generator Jacobian}}\Big{]}\end{split}$ (11)
-where we absorb $\frac{\partial\mathbf{\mathbf{\hat{z}}}_{t}}{\partial\mathbf{\mathbf{\hat{z}}}}$ and the other constant into $w(t)$.
+$\begin{split}&\nabla_{\theta}\mathcal{L}_{\mathrm{MSE}}(\boldsymbol{\hat{z}},\boldsymbol{z},c_{y})=\mathbb{E}_{t,\epsilon}\Big{[}w(t)\cdot\underbrace{\frac{\partial{\epsilon}_{\psi}(\boldsymbol{\hat{z}}_{t};t,c_{y})}{\partial\boldsymbol{\hat{z}}_{t}}}_{\text{Diffusion Jacobian}}\\ &\underbrace{({\epsilon}_{\psi}(\boldsymbol{\hat{z}}_{t};t,c_{y})-{\epsilon}_{\psi}(\boldsymbol{z}_{t};t,c_{y}))}_{\text{Prediction Residual}}\underbrace{\frac{\partial\boldsymbol{\hat{z}}}{\partial\theta}}_{\text{Generator Jacobian}}\Big{]}\end{split}$ (11)
+where we absorb $\frac{\partial\boldsymbol{\boldsymbol{\hat{z}}}_{t}}{\partial\boldsymbol{\boldsymbol{\hat{z}}}}$ and the other constant into $w(t)$.
 The computation of the Diffusion Jacobian term is computationally demanding, as it necessitates backpropagation through the Teacher Model.
@@
 11 by omitting Diffusion Jacobian:
-$\begin{split}&\nabla_{\theta}\mathcal{L}_{\mathrm{TSM}}(\mathbf{\hat{z}},\mathbf{z},c_{y})=\\ &\mathbb{E}_{t,\epsilon}\Big{[}w(t)\underbrace{({\epsilon}_{\psi}(\mathbf{\hat{z}}_{t};c_{y},t)-{\epsilon}_{\psi}(\mathbf{z}_{t};c_{y},t))}_{\text{Prediction Residual}}\underbrace{\frac{\partial\mathbf{\hat{z}}}{\partial\theta}}_{\text{Generator Jacobian}}\Big{]}\end{split}$ (12)
+$\begin{split}&\nabla_{\theta}\mathcal{L}_{\mathrm{TSM}}(\boldsymbol{\hat{z}},\boldsymbol{z},c_{y})=\\ &\mathbb{E}_{t,\epsilon}\Big{[}w(t)\underbrace{({\epsilon}_{\psi}(\boldsymbol{\hat{z}}_{t};c_{y},t)-{\epsilon}_{\psi}(\boldsymbol{z}_{t};c_{y},t))}_{\text{Prediction Residual}}\underbrace{\frac{\partial\boldsymbol{\hat{z}}}{\partial\theta}}_{\text{Generator Jacobian}}\Big{]}\end{split}$ (12)
 The effectiveness of the method can be proven by starting from the KL divergence.
-We can use a Sticking-the-Landing [39] style gradient by thinking of ${\epsilon}_{\psi}(\mathbf{z}_{t};c_{y},t)$ as a control variate for $\hat{\epsilon}$.
+We can use a Sticking-the-Landing [39] style gradient by thinking of ${\epsilon}_{\psi}(\boldsymbol{z}_{t};c_{y},t)$ as a control variate for $\hat{\epsilon}$.
 For detailed proof, refer to Appendix 4 of DreamFusion [36].
@@
 Specifically, we find that VSD employs identical inputs for both the Teacher and LoRA models to compute the gradient, while here TSM uses high-quality and suboptimal inputs for the Teacher Model.
-The losses are related to each other through $\epsilon_{\phi}(\mathbf{\hat{z}}_{t};t,c_{y})$.
+The losses are related to each other through $\epsilon_{\phi}(\boldsymbol{\hat{z}}_{t};t,c_{y})$.
 Appendix E Algorithm